A Complete Bibliography of Publications in Annals of Probability (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

13 October 2023 Version 1.04

Title word cross-reference

(2 + 1) [CD20b]. 2 [DG22, FL20, GG21]. 3 [AS20, GL21, HZZ23]. _{4} [ALS22]. b [Kry23]. β [DV23, FW21]. b ∈ L_{d} [Kry21b]. d ≥ 3 [SZ22]. Dσ [Kry23]. k [DGJ^{+}22]. κ ∈ (4, 8) [GP20]. L_{d} [Kry21a, Kry23]. Z^{2} [JRA20]. Z^{d} [Dem20]. SLE_{κ} [GP20]. N [SSZZ22]. o [FKS23]. O(N) [SSZZ22]. p [CDS23, FZK23]. σ ∈ W_{d}^{1} [Kry21b].

-circular [DV23]. -dimensional [CD20b]. -ensemble [FW21]. -minimal [FKS23]. -regular [DGJ^{+}22]. -sticking [CDS23]. -variation [FZK23].

1 [DL23]. 1-D [DL23].

2d [CSST21, IOSV22, Sch20b, CES21]. 2d-directed [CSST21].

Abelian [BR21]. above [CD22]. Absence [BDH23]. absolute [AP20, SV21]. added [HSS22]. Additive [DOP21, JNNP23]. adjacency [GP20]. adjacent [CLL20]. adsorbed [LP22]. Age [CRY21]. aggregation [GP21]. Airy [BGH21, DV21]. Aldous [LMW20]. algorithm [NW23]. Algorithmic [BGJ20]. algorithms [GJ21]. allocation [HPZ21]. Almost [BBPS22, CK20, DMX21, GHM20, HP20]. Almost-sure [BBPS22, CK20]. along [LP22]. among [BS20, DFSX21]. Anderson [ČD20a, CDOT21, CvZ21, DL23]. Anisotropic [Bla20, CES21]. annealed [Van21]. annihilating [AGJ21]. Anomalous [FT20]. Anti [FKS21]. Anti-concentration [FKS21]. any [AS21]. application [LSZ20]. applications [Aug20, MP22]. approach [KS20b, Sub23]. approaching [GL22]. approximate [DLS23, GJ21]. approximation [BCG20, CX22, JL21, RX21]. arbitrary [BDH23, SV21]. associated [ACHTS21]. asymmetric [BMRS20, Sel22]. Asymptotic [FTZ22, CK20]. Asymptotics [CvZ21, BMW20, MMT20]. Atlas [BB22]. attraction [BB22]. Atypical [BC21]. averages [CDSZ22]. Averaging [HL20]. avoiding [DCGHM20]. Aztec [Ber21].

backward [BDH23, DG22]. Balanced [AHR23, BCDG22]. ball [AJM21]. ballistic [CDS23]. Band [AEKS20]. Bang [HLYZ22]. Baxter [BM22b]. bees [BBNP22]. behavior [CK20, HLYZ22, LMS23]. Benamou [BVBHK20]. Bernoulli [CD22, DFSX21]. Berry [CDMP23]. beta [VV20]. beyond [BCDS21]. Big [HLYZ22]. binary [RST21]. binomial [MNPS20]. bipartite [BL22b]. bipolar [BM22b]. Board [Ano23a, Ano23b, Ano23c]. bootstrap [Bla20]. Boundary [FK23, HMP20]. bounded [HS20a]. Bounding [DCGHM20]. bounds [Aug20, CDMP23, EMZ20]. branching [AGJ21, BIM21, ČD20a, CRY21, HHKW22]. Brenier [BVBHK20]. Breuer [CNN20]. Brownian [AHS20, APPS22, BBNP22, CSST21, DM21, GOT22, GKP21, GHM20, HL20, HMP20, JNNP23, Jeg20, KS20a, KMP23, LR20a, LMN20, MS21, PR21, Rie22, SZ20, SV21]. bubbles [GP20]. Bulk [DV21]. Busemann [JRA20].

Capacity [Sch20a]. case [GOT22, LMS23]. causal [BCS22]. central [APPS22, DT23]. chain [LMW20]. chains [LMS23, MPS22]. chaos [BCD21, Jeg20, Lac22, NPS23, RZ20, SW20]. chaotic [CCL20]. Characterization [DM21, Eis14, Eis22]. charge [APPS22]. Chase [BCE^{+}21]. Chase-escape [BCE^{+}21]. cheese [AS20]. circular [DV23, FW21, VV20]. cladograms [LMW20]. class [FTZ22]. classes [Kry23]. CLE [ALS22]. closed [JL23, Sch22]. CLT [EMZ20]. clusters [DG21, GRS22]. coalescence [LR04, LR20b]. coalescent [BM22b]. coalescent-walk [BM22b]. coalescing [HLYZ22]. codings [Spi20a]. coefficients [BCDG22, HJ20, NPS23, XGL23]. collapsed [LP22]. collisions [CCL20]. Color [BB21b]. Color-position [BB21b]. colored [Gal21]. common [CDL20, HŠS21]. compact [Dem20, FT22]. Comparison [LLM20, HS20b]. complex [Lac22]. concave [CGM21, MP22]. concavity [CER23]. Concentration [MP22, DLR20, DGJ^{+}22, FKS21]. Concurrent [BGL23]. condensation [BMRS20]. condition [BCG20, SV21]. conditional [DT23]. conditioned [CN20, DFSX21]. conductances [BS20]. configuration [BCDS21]. Confluence [GM20]. Conformal [Lee21, ALS22, MSW22]. conjecture [CDD^{+}21, JSS21]. Connectivity [GP20]. consequences [EKK23]. constant [CD22]. constrained [MMT20]. constraints [DTV20, Nam20]. Constructing [CD20b]. contact [LMSV21]. containing [Kry23]. Contents [Ano23d, Ano23e, Ano23f]. continuity [AP20, MS21, SV21]. Continuous [CNN20, CD22, LW23]. continuous-time [LW23]. continuum [BS22]. control [DPT22]. Convergence [BR21, FL20, KMP23, Lac22, BB21a, HJ20, NW23]. converges [CSST21, GMS21]. convex [CER23, Dem20]. convolved [Bob23]. correction [LR20b]. Correlated [AEKS20, Eis14, Eis22]. correlation [BCG20]. correlations [CIM23, CGH21]. Coulomb [AS21]. counts [FKS21]. coupled [RX21]. coupling [BH22, KS20b, MV20]. covariance [BMR20, Lyo13, Lyo18, Lyo21]. cover [Sch20b]. Critical [IOSV22, ADK21, AOR21, BDW20, CKG23, DMX21, FT20, GG21, HHKW22, HMT21, Hut20, HMS23, LMS23, LMSV21, Pri21, Ros21, SW20]. criticality [GL22]. CRT [GMS21]. cumulants [FGR22]. current [Sch23]. curvature [CLW23]. curves [JL23]. Cut [HS21b, HH23]. Cut-off [HS21b]. Cutoff [CK22, HS20a, Sel22, HSS22, Sal23].

D [DL23, DG22, FL20, GG21, GL21, HZZ23]. damping [CX22]. Davies [JSS21]. death [BCE^{+}21]. Decay [Bob23]. degenerate [BS20]. degeneration [FK23]. degree [BNNS21]. degrees [BL22b, CKG23]. delay [KS20b]. delocalized [DL23]. densities [Bob23]. density [BMW20]. dependence [HHL20]. dependent [ABZ23, BR20, CX22, Spi20b]. depending [PT22a]. deposition [CDS23]. derivative [BIM21]. Derrida [CDD^{+}21]. determinantal [Més20]. Deterministic [AL20]. deviation [Aug20, PTV22, XGL23]. deviations [BGL23, BBG21, DLR20, GH20, Nam20]. diagonal [ACD^{+}21, KM20]. diamond [Ber21]. difference [BCDG22]. differential [ABM20, BCD21, GOT22, HJ20, RX21]. Diffusion [RX21, FT20, FK23, GP21, KOR20, LMW20, SS21, Sch22]. diffusion-limited [GP21]. Diffusions [FPRW21, DK20, KM20, LNS23]. digraphs [ABS23]. diluted [BCS23]. dimension [AS20, BDH23, DMX21, DCRRV23, GM21, Sch20a]. dimensional [ACHTS21, BS22, CSZ20, CX22, CD20b, FT20, HMS23, Sch20c, Tol21]. dimensions [AHR23, Bla20, CCL20, CvZ21, EMZ20, Lam21]. dimer [Lam21]. Dimers [BLR20]. directed [BC20a, CSST21, Dau22, DSV22, DCKN^{+}20, JRA20]. directions [HMT21]. Dirichlet [FPRW21, Sch22]. disconnected [PT22a]. disconnection [CN20]. discontinuities [Ben20]. Discrete [MPS22, Ger20]. disordered [BMRS20, BC20b, GL22]. displacement [GMMZ22, GH22, GM21].

Distance [Lyo13, Lyo18, Lyo21, ALS22]. Distribution [DFSX21, BR20, BC20a, Liu22, RZ20]. distributional [ABM20, Lee21]. distributions [BB22, CGM21, Dia23, FPRW21, MP22, SZ22]. divisible [BM22a]. domain [Tol21]. Domains [BB22]. Domino [Ber21]. Donsker [BGL23]. doubly [Ber21]. DPRZ [Sch20b]. drift [ABM20, Kry21a, LMN20]. drifted [SZ20]. driven [GOT22, HL20]. duality [BL22a]. Duarte [MMT20]. during [HLYZ22]. dynamic [DPT22]. dynamical [BHT20, BZG23, HS20b, Ros21]. dynamics [HL20].

edge [AEKS20, ADK23, BBG21]. edges [CH20]. Editorial [Ano23a, Ano23b, Ano23c]. eigenvalue [GH20, LLM20]. eigenvalues [ADK21, CvZ21, Dia23, HLY20]. Eigenvector [ALM21]. Elliptic [ADG20, BCDG22, DG21]. embedding [EMZ20, GMS21, MS21]. Emergence [CS21]. endpoint [BC20a]. energy [BCS23, FMM21, KM20, Sub23]. ensemble [DV21, FW21, VV20]. ensembles [DM21, MSW22, Nam20]. enstrophy [FL20]. entire [CSZ20]. Entrance [DK20]. Entropic [CN20]. entropy [KS23, Més20]. environment [AL20, ČD20a, JM20, dR22]. environments [BHT20]. epidemics [BNNS21]. equation [AJM21, BC23, CSZ20, CD20b, CK20, CK23, CGH21, DG23, DLR20, DG22, KOR20, NQR20, Tol21]. equations [ABM20, BCD21, BR20, BBPS22, BCDG22, BGW22, DMX21, FL20, FT20, FK23, GMMZ22, GOT22, HZZ23, HJ20, Kry21a, Kry21b, Kry23, KS20b, Pri21, RX21, SS21]. equilibrium [BC21, BB21a, LNS23]. Erdos [ADK21, ADK23, Aug20, HLY20, Ros21]. Ergodic [JRdlR20, MPS22]. ergodicity [CES22, DKM20]. Errata [CDL20, Lyo18, Lyo21]. Erratum [Eis22]. escape [BCE^{+}21]. Esseen [CDMP23]. estimates [CDOT21, CLW23, FZK23, GOT22, KM20, PTV22]. Euclidean [Tim21]. Euler [FL20]. events [BC21, BMR20]. evolution [CRY21]. Exact [MMT20]. example [BZG23]. Exchangeable [Ger20]. excited [AHR23]. exclusion [HP20, JM20, Sal23]. excursion [GRS22]. Existence [DCRRV23, HZZ23, ABM20, HŠS21]. exit [BC21, DK20, SS21]. expanders [ABS23]. Expansion [EKK23]. expansions [FTZ22, XGL23]. Exponential [BZG23, BBPS22, Hut20]. exponents [LNS23, LMSV21]. extended [CS21]. External [GP21]. extinction [BNNS21]. extrema [KMP23]. Extremal [ADK21, ALS22, HLY20, LLM20]. extreme [PT22b].

face [BL22b]. faces [BCS22]. factor [Tim21]. factors [SZ22]. faster [HP20]. Ferguson [Sch22]. field [BCD21, BH22, BS22, CDL20, CN20, CRY21, DLR20, EMS21, GMMZ22, GL22, GRS22, HS20a, HLYZ22, NS20]. fields [BMR20, BW20, DW20, DCRRV23, Spi20a]. Finitary [KS20a, Spi20a, Spi20b]. Finite [HMT21, HHL20]. Finitely [Spi20b]. fire [CRY21]. first [BDH23, Dem20]. first-passage [BDH23]. fixate [AGJ21]. fixed [SV21]. FK [Izy22]. FKG [MRVKS23]. flow [Dem20, LMN20]. Flows [LR04, LR20b]. fluctuation [CES22, GH22, LNS23]. fluctuations [HLY20, MP22]. Fokker [BR20, FT20]. Forest [PT22a, CSST21, CRY21]. Forests [FGR22, BDW20]. formula [BMR20, Ben20, MV20]. forward [DG22]. forward-backward [DG22]. four [Sch20c]. four-dimensional [Sch20c]. Fractal [BGH21, DG23]. Fractional [KOR20, GOT22, HL20, JNNP23]. fragment [DGJ^{+}22]. fragmentation [DGJ^{+}22, LR20a, Thé21]. Fredholm [FTZ22]. Free [BCS23, PT22a, DW20, FMM21, GL22, GRS22, Sub23]. Freeness [ACD^{+}21]. Frozen [RST21]. full [KM20]. fully [RX21]. function [AOR21, Pim22, SW20]. functionals [DOP21, JNNP23]. functions [DG21, JRA20].

Galton [dR22]. game [DLR20]. games [CDL20, GMMZ22]. gamma [BC23]. gap [GJ21]. gaps [CCL20, FW21]. gas [DV23]. gases [AS21]. Gaussian [Eis22, BC20b, BMR20, Ben20, DW20, DCRRV23, Eis14, FFN21, GRS22, HLY20, Jeg20, Lac22, MRVKS23, RZ20, SW20, Tho23]. General [Rao20, BNNS21, XGL23]. generalisation [Lam21]. generalized [KS20b]. generator [PT22a]. genus [BL22b, JL23]. Geodesic [Le 22, BF22]. Geodesics [MS21, DSV22, GM20]. Geometric [DKM20, FKS23, CH20, Thé21]. geometry [BGH21, BLR20, GG21, Le 22, Lee21]. Gibbs [JRA20, Rao20]. Gibbsian [DM21]. Ginzburg [BW20]. given [Nam20]. glass [BCS23]. glasses [EMS21, FMM21, Sub23]. Global [BTZ23, HZZ23, Tol21, BPW21, GG21]. Global-in-time [HZZ23]. Gordon [BH22]. graph [BTZ23, CKG23, GP20]. Graphexes [BCDS21]. graphs [ADK21, ADK23, Aug20, BNNS21, BBG21, CH20, CIM23, CK22, CS21, DW20, FKS21, HSS22, HS21b, HLY20, Hut20, Lee21, LMSV21, Ros21]. Gravitational [HPZ21]. gravity [GM20, GMS21, MS21, MSW22]. Green [DG21]. group [XGL23]. groups [FT22, PT22a]. Growth [LR20a, Hut20, JR22, Lee21]. Growth-fragmentation [LR20a].

half [BC23, BL22a]. half-plane [BL22a]. half-space [BC23]. halves [DSV22]. Hamiltonian [BCS23, CvZ21, DL23]. Hamiltonians [GMMZ22]. Hammersley [DCGHM20]. hard [CCL20, LP22]. Harnack [BCDG22]. heat [AJM21, CLW23, CK20, CK23, DMX21, DG22, KM20]. heavy [CDS23]. height [ABC21]. Hermitian [DP21]. Hidden [Dau22, PT22b]. High [HMS23, BL22b, EMZ20]. High-dimensional [HMS23]. higher [Lam21]. Hitting [LMN20, SZ20, NS20]. holomorphic [NPS23]. homogenization [CCKW21, DG21, FK23, GGM22]. honeycomb [Lam21]. hydrodynamical [BGL23]. Hydrodynamics [BMRS20]. hyperbolic [JL23, LMSV21]. Hypercontractivity [PTV22]. hypercube [EKK23]. hypergraphs [Ger20]. hypersurface [DCRRV23]. Hypoelliptic [GOT22].

i.i.d [CKG23]. Identification [MV20]. II [MS21, Sub23]. IID [SZ22]. III [Lac22]. imaginary [BLR20]. Improved [CER23]. independent [HP20]. indexed [LR20a]. inequalities [CGM21, MP22, Rie22]. inequality [BCDG22]. inference [FMM21]. infinite [AHS20, BB22, BBNP22, BDH23, BL22a, CX22]. infinitely [BM22a, HH23]. infinity [Dem20, DK20]. information [BTZ23]. inhomogeneous [JR22, LTV21]. initial [SV21]. insertion [DCGHM20]. Integration [Pim22]. interacting [BB21b, FTZ22, GGM22, KMP23, LNS23, LP22, SZ20]. interaction [NS20]. interface [GG21, KOR20]. interfaces [GL21]. interlacements [CN20]. intermittency [CK23]. interval [FPRW21, Ger20, RZ20]. intervals [AOR21]. Invariance [JM20, BS20, ČD20a, Dau22]. invariant [BB22, CCL20, CER23, Rao20]. Inverting [LSZ20]. Ising [BS22, CIM23, GG21, GL21, IOSV22, Izy22, Rao20, Xu21]. islands [CK23]. isomorphisms [KS20a]. Isoperimetric [Rie22]. isoradial [CIM23]. Itô [Kry21b, Kry23]. Itô’s [Ben20]. iterated [DG23].

jump [DK20, SZ20].

Kac [CCL20]. KCM [HMT21]. kernel [CLW23, KM20]. kinetic [FT20, KOR20]. kinetically [MMT20]. Kolmogorov [NQR20, Pri21]. KPZ [BC23, CES21, CSZ20, CD20b, CGH21, DG23, GHM20, LNS23, Pim22, SV21]. KPZ-type [LNS23]. Kramers [CX22].

Landau [BW20]. landscape [CK23, DSV22]. Laplace [Bob23, PT22b]. Large [DLR20, GH20, Nam20, SSZZ22, XGL23, ACD^{+}21, Aug20, BGL23, BBG21, BCS22, JL23]. largest [DGJ^{+}22, GH20]. last [BF22, Dau22, DCKN^{+}20]. last-passage [DCKN^{+}20]. lattice [GL22]. Law [DG23, CKG23, Lac22]. laws [AS21, PT22b]. least [KOV20]. left [CDMP23]. level [DW20]. level-sets [DW20]. Lévy [ALM21, CCKW21, CK20, CK23]. Lévy-type [CCKW21]. lifts [CK22]. Limit [JNNP23, Tey20, BBNP22, CDS23, CKG23, DLR20, DT23, HHKW22, KOR20, LMW20, PT22b, Ros21, SSZZ22, VV20, dR22]. limited [GP21]. Limiting [Més20, Dia23]. Limits [BCDS21, ACHTS21, BM22b, BL22b, Ger20, Lee21, Tho23]. line [AOR21, DV21, DM21, SW20]. linear [SSZZ22, XGL23]. Liouville [APPS22, GM20, GMS21, MS21, MSW22]. Lipschitz [CD22]. Littlewood [FKS23]. Local [AS21, BL22b, GG21, AHS20, BTZ23, BM22b, GOT22, LSZ20]. Locality [ABS23, Hut20]. Localization [BC20b, CH20, ADK23, Nam20]. locally [FT22, Sjö23]. log [BC23, CER23, CGM21, MP22]. log-concave [CGM21, MP22]. log-concavity [CER23]. log-gamma [BC23]. log-Sobolev [CGM21]. Logarithmic [CLW23]. logarithms [DG23]. loop [ALS22, MSW22]. loops [APPS22]. low [BC20b]. Lower [KS23, BBG21, PTV22].

Major [CNN20]. Mallows [ABC21, HHL20]. manifold [Sch22]. many [CH20, HH23]. map [BL22a, GP21, GMS21, MS21]. maps [BCS22, BL22b, GM21, JL23]. Markov [CCL20, HHKW22, HZZ23, LMS23, MPS22, Spi20a]. Markovian [LSZ20]. Martingale [BVBHK20, BIM21, EMZ20, FZK23]. martingales [FGR22, GKP21]. master [DLR20, GMMZ22]. matching [GH22, HSS22]. mated [GMS21]. mated-CRT [GMS21]. matrices [ALM21, AEKS20, ACD^{+}21, Aug20, CES22, Dia23, DP21, DLS23, GH20, JL21, LTV21]. matrix [KOV20]. maximal [ČD20a, Dem20, DCKN^{+}20, Sch23]. Maximum [BH22, BW20, GL21, Sch20c]. McKean [DPT22, HŠS21]. Mean [GMMZ22, HLYZ22, NS20, BCD21, CRY21, DLR20, EMS21, HS20a, CDL20]. mean-field [EMS21, HS20a]. measure [DLR20, FL20, MV20]. Measures [AHS20, BC23, CCL20, CER23, JRA20, KM20, Sch22]. medium [BGW22]. membrane [Sch20c, Tho23]. message [DLS23, GJ21]. Metastability [LMSV21, SS21, FK23]. Metastable [LMS23]. method [BMW20, DT23]. metric [CKG23, DW20, Lyo13, Lyo18, Lyo21]. microcanonical [Nam20]. minimal [FKS23]. mixes [HP20]. Mixing [CLL20, Sch23, BBPS22, BHT20, BZG23, LMS23, Spi20a]. model [BB22, BTZ23, BCS23, BS22, ČD20a, CIM23, CDS23, CRY21, GL22, HS21a, IOSV22, Izy22, KP22, Lam21, Rao20, Sch20c, SSZZ22, SZ22, Tho23, Xu21]. models [BCDS21, CDD^{+}21, CDOT21, Gal21, JRA20, LSZ20, MMT20, MRVKS23, Sub23]. Modified [CGM21]. Moment [CDOT21, DTV20]. Moments [AOR21]. Monge [GKP21]. Monotone [Sjö23, HS20a, HJ20]. Monotonicity [LW23, GMMZ22]. Morrey [Kry23]. Most [HH23]. motion [AHS20, GOT22, GHM20, HL20, HMP20, JNNP23, Jeg20, KMP23, LR20a, PR21]. motions [KS20a, SZ20]. multi [FT20]. multi-dimensional [FT20]. MultiGraphexes [BCDS21]. multinomial [KP22]. multiple [BPW21, Izy22, Nam20]. multiplicative [Jeg20, Lac22, NPS23, RZ20, SW20]. multiplicity [AHS20]. Multipoint [Liu22]. multispecies [Sub23]. multitype [CRY21]. Multivariate [JL21].

nature [AS20]. Navier [BBPS22, HZZ23]. Near [BDW20, BC21, HMS23]. Near-critical [BDW20, HMS23]. negatively [ABZ23]. networks [NS20]. nodal [AP20, DCRRV23]. noise [CK20, CK23, CvZ21, FL20, HŠS21, LR04, LR20b, CDL20]. noisy [BTZ23]. non [DP21, Tho23]. non-Gaussian [Tho23]. non-Hermitian [DP21]. nonamenable [Tim21]. nonconvergence [KMP23]. nonendogenous [RST21]. nonexistence [MNPS20]. nonglobally [HJ20]. Nonlinear [Aug20, BR20, DMX21, DG22, Tol21]. nonlocal [HHKW22]. nonseparable [GMMZ22]. nonstationarity [CNN20]. nonsymmetric [CCKW21]. nontriviality [CES21]. nonuniform [BHT20]. nonuniqueness [HZZ23]. Normal [BCG20, CES22, JL21]. normed [PTV22]. Number [DV23, DCGHM20, DCKN^{+}20, HMT21]. Number-rigidity [DV23]. observations [BTZ23]. obstacles [DFSX21]. occupation [CN20]. occupation-time [CN20]. off [HS21b, KM20]. off-diagonal [KM20]. Offord [FKS23]. Operator [VV20]. Optimal [BGW22, DPT22, GH22, Pri21]. Optimization [EMS21]. order [BCG20, PT22b]. ordered [Ger20]. ordinary [HJ20]. orientations [BM22b]. Ornstein [FL20]. orthant [HS21a]. overlap [GJ21].

parabolic [ČD20a, CDOT21, DG21, FK23]. partial [HJ20]. particle [BB21b, ČD20a, FTZ22, GGM22]. particles [HP20]. partitions [FPRW21]. parts [Pim22]. passage [BDH23, BF22, Dau22, Dem20, DCKN^{+}20]. passing [DLS23, GJ21]. passive [BBPS22]. paths [BDH23, DOP21, DCKN^{+}20]. PCA [BGJ20]. peaks [CK23]. Pekar [MV20]. perceptron [Xu21]. Percolation [DW20, Bla20, BDH23, BF22, CD22, DG21, Dau22, Dem20, DCKN^{+}20, HS20b, HMS23, MRVKS23, RST21, Ros21, Van21]. Periodic [BMW20, CCKW21, Ber21]. permutations [BM22b, HHL20, Sjö23]. Persistence [FFN21]. perspective [BVBHK20, FFN21]. perturbation [HJ20]. Perturbations [FK23]. perturbed [KMP23]. Pfaffians [FTZ22]. phase [CDSZ22, DL23, Lac22, LP22, MRVKS23, Sch23]. Pierrehumbert [BZG23]. pinning [GL22]. Planar [Jeg20, AHS20, GP21, GM21, MRVKS23]. Planck [BR20, FT20]. plane [BL22a, Rie22]. plateau [HMS23]. point [Pim22, SV21]. Points [AHS20, DMX21, HPZ21]. Poisson [ADK23, FPRW21, JRdlR20]. Polarity [DMX21]. Polaron [MV20]. Pólya [BMW20]. polygon [DCGHM20]. polymer [BC23, JRA20]. polymers [BC20a, Dau22]. polynuclear [JR22]. porous [BGW22]. posedness [KS20b, Tol21]. position [BB21b]. positively [Eis14, Eis22]. potential [CvZ21]. potentials [NW23]. potlatch [BB21a]. power [CKG23]. power-law [CKG23]. Precise [GOT22]. prescribed [BL22b]. prewetting [GG21, IOSV22]. principle [BS20, DPT22, HS20b, JM20]. principles [ČD20a]. probabilistic [BVBHK20]. probabilistically [HZZ23]. probabilities [AJM21, LMN20]. probability [Hut20, MNPS20, Sch22]. problem [GKP21, GH22]. problems [DTV20, FKS23, FK23, SS21]. process [BMRS20, CCL20, HS20a, HP20, LMSV21, SZ20]. processes [ABZ23, BB21a, BM22a, Ben20, BM22b, CCL20, CCKW21, CRY21, Eis14, Eis22, FFN21, FK23, HHKW22, LMS23, LR20a, Més20, Spi20b, Thé21]. products [KOV20]. profile [Tey20]. programming [DPT22]. projection [LSZ20]. proof [Sch20b]. Propagation [BCD21]. properties [DG23, DV21, GP20]. property [GJ21]. pure [Sub23].

quadratic [BCS23]. Quantitative [DG21, EMZ20, GGM22]. quantization [ADG20, SSZZ22]. quantum [CES22, GM20, GMS21, MS21, MSW22]. Quenched [BS20, ČD20a, DFSX21].

Rademacher [GH20]. radial [LMN20]. radius [ALS22, GRS22]. Random [BHT20, DTV20, GM21, KOV20, AGJ21, AL20, ABS23, AEKS20, APPS22, ACHTS21, AHR23, AP20, ACD^{+}21, BS20, BCDG22, BNNS21, BBG21, BCS22, BZG23, BR21, BS22, BIM21, ČD20a, CH20, CN20, CK22, CKG23, CS21, CDMP23, DFSX21, DP21, EKK23, FKS21, GP21, HH23, HS20b, HLYZ22, HSS22, JM20, JL21, KMP23, Le 22, LMSV21, LTV21, LW23, MP22, PT22b, Ros21, Sjö23, Spi20a, Tey20, XGL23, dR22]. range [BMRS20, HS20a, HMP20, LMS23, Sch20a]. Rates [BB21a, HS20a, HJ20, Lee21]. reaction [SS21]. real [JSS21]. Recurrence [BL22a]. recursive [CDD^{+}21]. regime [BMRS20, CSZ20, HLYZ22]. regular [DGJ^{+}22]. regularity [BGW22, Pri21]. reinforced [SZ20]. relation [GHM20]. relative [KS23]. renormalization [BDW20]. renormalized [CDOT21]. Rényi [ADK21, ADK23, Aug20, HLY20, Ros21]. repeated [CDSZ22]. repelling [BC21]. representation [Thé21]. repulsion [CN20]. reservoirs [Sal23]. resistance [GM21]. resistances [BL22a]. restrictions [CLW23]. result [GH22, Pri21]. Retaux [CDD^{+}21]. reversible [CCL20, LMS23]. Riemann [AOR21, SW20]. Riemannian [Sch22]. Riesz [DV23]. riffle [Sel22]. rigidity [AEKS20, AS21, DV23]. rooted [MMT20]. rotationally [CER23]. Rough [FZK23, BCD21, DOP21, PR21].

sample [Van21]. sandpile [BR21]. sandpiles [HS21b]. scalars [BBPS22]. scaled [KMP23]. Scaling [ACHTS21, BM22b, CES21, CDS23, Ros21, dR22, CKG23, Tho23]. Schrödinger [NW23]. SDE [BR20, DG22]. SDEs [HŠS21]. Second [Lyo21, BCG20]. second-order [BCG20]. Secular [NPS23]. self [DCGHM20, DGJ^{+}22, KMP23, LP22, MPS22]. self-avoiding [DCGHM20]. self-interacting [KMP23, LP22]. self-similar [DGJ^{+}22, MPS22]. semimartingales [FZK23]. semirandom [DLS23]. sensitivities [KS20b]. set [PT22a]. sets [CER23, DW20]. setting [Rao20]. Shannon [FT22]. shape [HS21a]. Sharp [DGJ^{+}22, Xu21]. short [AOR21]. shredded [BCS22]. shuffle [Sel22]. sigma [SSZZ22]. similar [DGJ^{+}22, MPS22]. Simple [JL23, MSW22, Sch20b]. simplex [CLL20]. sine [BH22]. sine-Gordon [BH22]. singular [KOV20, LTV21, NS20]. singularity [KM20]. Sinkhorn [NW23]. SLE [GHM20, Izy22]. SLEs [BPW21]. Slicing [CNT22]. Small [AJM21, FW21]. smallest [DGJ^{+}22, LTV21]. Smoluchowski [CX22]. smooth [BMR20, DCRRV23]. smoothing [BB21a]. Sobolev [CGM21, DKM20]. Solution [NQR20, CD20b, CK20, GKP21]. solutions [BR20, HZZ23, Kry21b, Kry23]. some [CDOT21, LLM20, PT22a, Pri21]. space [BC23, BGW22, CKG23, DKM20, FPRW21, Sch22, Tim21]. spaces [Lyo13, Lyo18, Lyo21, PTV22]. Spanning [PT22a, ACHTS21, BDW20, CSST21, GP21]. spanning-tree-weighted [GP21]. sparse [ADK23, Aug20, BBG21, BCDS21, CS21, HLY20]. spatial [Spi20a]. SPDEs [Pri21]. Spectral [BBG21, CCL20, GM21, ADK23, FFN21, Lee21, Van21]. spectrum [CS21]. sphere [CCL20, HPZ21]. spheres [BCS22]. spherical [Sub23]. spin [BCS23, CIM23, EMS21, FMM21, Sub23]. splittings [JRdlR20]. square [LTV21]. squared [Eis14, Eis22]. Stability [NW23, KS20b]. Stable [BCS22, ABM20, CKG23, DK20, HMT21, Thé21]. stars [Le 22]. state [CX22]. state-dependent [CX22]. states [CS21, Rao20]. stationarity [CES21]. Stationary [BC23, SZ22, FFN21, FPRW21]. Statistics [SW20, ALM21, ADK23, LLM20, PT22b]. Stein [DT23]. sticking [CDS23]. stochastic [ADG20, ABZ23, ABM20, AJM21, BBPS22, Ben20, BGW22, CK20, CK23, DMX21, DG22, Gal21, HZZ23, HJ20, Kry21a, KS20b, LSZ20, RX21, SS21, SSZZ22, Tol21]. Stokes [BBPS22, HZZ23]. strictly [CD22]. Strong [ABM20, HZZ23, HJ20, Kry21b, Kry23, MV20]. strongly [CGM21]. structures [Ger20]. subcritical [CSZ20]. subdiffusive [dR22]. subgraph [FKS21]. subgraphs [EKK23]. subsequences [Sjö23]. subset [Dem20]. subsets [MNPS20]. sums [BCG20]. super [HMP20, PR21]. super-Brownian [HMP20, PR21]. supercritical [EKK23, MMT20]. support [AJM21]. suprema [BM22a]. sure [BBPS22, CK20, GHM20]. Surface [LP22, APPS22, MP22]. surfaces [JL23]. Survival [BNNS21, DFSX21]. swarm [BBNP22]. Swiss [AS20]. Symmetric [JM20, CER23, KM20]. Symmetries [Gal21, PT22b]. symmetry [BB21b]. system [CX22]. systems [BC20b, BZG23, BB21b, DMX21, FTZ22, GGM22, NS20, SS21].

Table [Ano23d, Ano23e, Ano23f]. tableaux [BMW20]. tail [ABZ23, BBG21]. tailed [CDS23]. tails [GL21, KS23]. TAP [FMM21, Sub23]. TASEP [Liu22, NQR20, Sch23]. temperature [AS21, BC20b]. tensor [BGJ20]. theorem [AJM21, CNN20, DT23, FT22, HS21a, LLM20, Sch20b]. theorems [JNNP23]. theory [DLR20, HJ20]. Thermodynamic [Tho23]. Three [DSV22, ACHTS21, Bla20, CCL20]. three-dimensional [ACHTS21]. Three-halves [DSV22]. threshold [Xu21]. thresholds [BGJ20]. Tightness [CNN20, GL21, CES21]. tiling [HS21b, KP22]. tilings [Ber21]. time [BGW22, CLL20, CD22, CN20, CGH21, HZZ23, LW23]. times [AHS20, HH23, NS20, SZ20, Sch23, Sch20b]. too [CH20]. topological [BMR20, HMP20]. torus [HMS23]. traces [JL21]. Tracy [HLY20]. transform [Bob23]. transforms [FZK23]. transient [HH23]. Transition [HLY20, CDSZ22, LP22, MRVKS23]. transitive [Hut20]. Translation [Rao20]. Translation-invariant [Rao20]. transport [FL20, GKP21]. transpositions [Tey20]. tree [ACHTS21, BF22, GP21, LR20a, RST21, dR22]. trees [ABC21, BCE^{+}21, Thé21]. triviality [ABZ23]. Tutte [GMS21]. two [AGJ21, AHR23, BS22, CSZ20, CvZ21, Pim22, Tol21]. two-dimensional [BS22, CSZ20, Tol21]. two-point [Pim22]. two-type [AGJ21]. type [AGJ21, CCKW21, CDMP23, LNS23].

Uhlenbeck [FL20]. unbounded [DCRRV23, Tol21]. Unicellular [JL23]. Uniform [PT22a, ACHTS21, BL22a, HPZ21, Sjö23]. uniqueness [ABM20, BPW21, HŠS21, Pri21]. unit [RZ20]. unitary [JL21]. Universality [BF22, CIM23, DLS23, HMT21, HSS22, KOV20, Sal23, AEKS20]. Upper [BBG21]. urns [BMW20].

value [LTV21]. values [KOV20]. Varadhan [BGL23]. variation [DSV22, FZK23]. variational [FMM21, MV20]. vertex [Gal21, SZ20]. via [Bob23, BL22a, CRY21, EMZ20, KS23, SSZZ22]. virtually [PT22a]. Vlasov [DPT22, HŠS21]. volatility [LSZ20]. volumes [AP20]. Voronoi [Van21]. voter [BTZ23, SZ22]. vs [JL23].

walk [ACHTS21, AHR23, BM22b, BIM21, CLL20, ČD20a, CDMP23, DFSX21, GM21, HS20b, LP22, PT22b, dR22]. walks [AGJ21, AL20, BS20, BHT20, DCGHM20, HH23, HLYZ22, KMP23, LW23, XGL23]. wall [LP22]. Watson [dR22]. wave [Tol21]. Weak [HŠS21, Pri21]. weakly [LMS23]. web [CSST21]. weighted [BCG20, CK22, DKM20, GP21]. weightings [Ber21]. Well [KS20b, Tol21]. Well-posedness [KS20b]. Welsh [DCGHM20]. white [CvZ21]. Widom [HLY20]. Wigner [Aug20, CES22, Dia23]. without [CLW23, MRVKS23]. words [DP21].

Yaglom [HHKW22]. Young [BMW20].

zero [BMRS20, CS21, HS20a, LMS23]. zero-range [BMRS20, HS20a, LMS23]. zeta [AOR21, SW20].

References

Addario-Berry:2021:HMT

[ABC21] Louigi Addario-Berry and Benoît Corsini. The height of Mallows trees. Annals of Probability, 49(5):2220–2271, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

           (electronic). URL https://projecteuclid.org/journals/
           annals-of-probability/volume-49/issue-5/The-height-of-
           Mallows-trees/10.1214/20-AOP1503.full.

Athreya:2020:SEU

[ABM20] Siva Athreya, Oleg Butkovsky, and Leonid Mytnik. Strong existence and uniqueness for stable stochastic differential equations with distributional drift. Annals of Probability, 48(1):178–210, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

           894X (electronic). URL https://projecteuclid.org/euclid.
           aop/1585123326.

Alimohammadi:2023:LRD

[ABS23] Yeganeh Alimohammadi, Christian Borgs, and Amin Saberi. Locality of random digraphs on expanders. Annals of Probability, 51(4):1249–1297, July 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-4/
Locality-of-random-digraphs-on-expanders/10.1214/22-AOP1618.
full.

Alishahi:2023:TTN

[ABZ23] Kasra Alishahi, Milad Barzegar, and Mohammadsadegh Zamani. On tail triviality of negatively dependent stochastic processes. Annals of Probability, 51(4):1548–1558, July 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-4/On-tail-triviality-of-
negatively-dependent-stochastic-processes/10.1214/23-
AOP1626.full.

Au:2021:FDL

[ACD^{+}21] Benson Au, Guillaume Cébron, Antoine Dahlqvist, Franck Gabriel, and Camille Male. Freeness over the diagonal for large random matrices. Annals of Probability, 49(1):157– 179, January 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-1/
Freeness-over-the-diagonal-for-large-random-matrices/
10.1214/20-AOP1447.full.

Angel:2021:SLT

[ACHTS21] O. Angel, D. A. Croydon, S. Hernandez-Torres, and D. Shiraishi. Scaling limits of the three-dimensional uniform spanning tree and associated random walk. Annals of Probability, 49(6): 3032–3105, November 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-6/
Scaling-limits-of-the-three-dimensional-uniform-spanning-
tree-and/10.1214/21-AOP1523.full.

Albeverio:2020:ESQ

[ADG20] Sergio Albeverio, Francesco C. De Vecchi, and Massimiliano Gubinelli. Elliptic stochastic quantization. Annals of Probability, 48 (4):1693–1741, July 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-4/
Elliptic-stochastic-quantization/10.1214/19-AOP1404.full.

Alt:2021:EEC

[ADK21] Johannes Alt, Raphaël Ducatez, and Antti Knowles. Extremal eigenvalues of critical Erdős–Rényi graphs. Annals of Probability, 49(3):1347–1401, May 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-3/
Extremal-eigenvalues-of-critical-Erd%c5%91sR%c3%a9nyi-
graphs/10.1214/20-AOP1483.full.

Alt:2023:PSL

[ADK23] Johannes Alt, Raphael Ducatez, and Antti Knowles. Poisson statistics and localization at the spectral edge of sparse Erdős– Rényi graphs. Annals of Probability, 51(1):277–358, January 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-1/Poisson-statistics-and-
localization-at-the-spectral-edge-of-sparse/10.1214/22-
AOP1596.full.

Alt:2020:CRM

[AEKS20] Johannes Alt, László Erdős, Torben Krüger, and Dominik Schröder. Correlated random matrices: Band rigidity and edge universality. Annals of Probability, 48(2):963–1001, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/euclid.aop/
1587542684.

Ahlberg:2021:FFT

[AGJ21] Daniel Ahlberg, Simon Griffiths, and Svante Janson. To fixate or not to fixate in two-type annihilating branching random walks. Annals of Probability, 49(5):2637–2667, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-5/To-fixate-or-not-to-
fixate-in-two-type-annihilating/10.1214/21-AOP1521.full.

Angel:2023:BER

[AHR23] Omer Angel, Mark Holmes, and Alejandro Ramirez. Balanced excited random walk in two dimensions. Annals of Probability, 51 (4):1421–1448, July 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-4/
Balanced-excited-random-walk-in-two-dimensions/10.1214/
23-AOP1622.full.

Aidekon:2020:PIM

[AHS20] Elie Aïdékon, Yueyun Hu, and Zhan Shi. Points of infinite multiplicity of planar Brownian motion: Measures and local times. Annals of Probability, 48(4):1785–1825, July 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-4/Points-of-infinite-multiplicity-
of-planar-Brownian-motion--Measures/10.1214/19-AOP1407.
full.

Athreya:2021:SBP

[AJM21] Siva Athreya, Mathew Joseph, and Carl Mueller. Small ball probabilities and a support theorem for the stochastic heat equation. Annals of Probability, 49(5):2548–2572, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-5/Small-ball-probabilities-
and-a-support-theorem-for-the-stochastic/10.1214/21-AOP1515.
full.

Aimino:2020:DWR

[AL20] Romain Aimino and Carlangelo Liverani. Deterministic walks in random environment. Annals of Probability, 48(5):2212– 2257, September 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-5/
Deterministic-walks-in-random-environment/10.1214/19-
AOP1421.full.

Aggarwal:2021:ESL

[ALM21] Amol Aggarwal, Patrick Lopatto, and Jake Marcinek. Eigenvector statistics of Lévy matrices. Annals of Probability, 49 (4):1778–1846, July 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-4/
Eigenvector-statistics-of-L%c3%a9vy-matrices/10.1214/
20-AOP1493.full.

Aru:2022:EDC

[ALS22] Juhan Aru, Titus Lupu, and Avelio Sepúlveda. Extremal distance and conformal radius of a CLE_{4} loop. Annals of Probability, 50 (2):509–558, March 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-2/
Extremal-distance-and-conformal-radius-of-a-CLE4-loop/
10.1214/21-AOP1538.full.

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial board. Annals of Probability, 51 (3):??, May 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-3/
Editorial-Board/aop513p1.full.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial board. Annals of Probability, 51 (4):??, July 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-4/
Editorial-Board/aop514p1.full.

Anonymous:2023:EBc

[Ano23c] Anonymous. Editorial board. Annals of Probability, 51(5): ??, September 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-5/
Editorial-Board/aop515p1.full.

Anonymous:2023:TCa

[Ano23d] Anonymous. Table of contents. Annals of Probability, 51 (3):??, May 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-3/
Table-of-Contents/aop513c1.full.

Anonymous:2023:TCb

[Ano23e] Anonymous. Table of contents. Annals of Probability, 51 (4):??, July 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-4/
Table-of-Contents/aop514c1.full.

Anonymous:2023:TCc

[Ano23f] Anonymous. Table of contents. Annals of Probability, 51 (5):??, September 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-5/
Table-of-Contents/aop515c1.full.

Arguin:2021:MRZ

[AOR21] Louis-Pierre Arguin, Frédéric Ouimet, and Maksym Radziwiłł. Moments of the Riemann zeta function on short intervals of the critical line. Annals of Probability, 49(6):3106–3141, November 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-6/Moments-of-the-
Riemann-zeta-function-on-short-intervals-of/10.1214/21-
AOP1524.full.

Angst:2020:ACR

[AP20] Jürgen Angst and Guillaume Poly. On the absolute continuity of random nodal volumes. Annals of Probability, 48(5):2145– 2175, September 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-5/
On-the-absolute-continuity-of-random-nodal-volumes/10.
1214/19-AOP1418.full.

Ang:2022:BLC

[APPS22] Morris Ang, Minjae Park, Joshua Pfeffer, and Scott Sheffield. Brownian loops and the central charge of a Liouville random surface. Annals of Probability, 50(4):1322–1358, July 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-4/Brownian-loops-and-the-
central-charge-of-a-Liouville-random/10.1214/21-AOP1558.
full.

Asselah:2020:NSC

[AS20] Amine Asselah and Bruno Schapira. On the nature of the Swiss cheese in dimension 3. Annals of Probability, 48(2):1002–1013, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL https://projecteuclid.org/euclid.
aop/1587542685.

Armstrong:2021:LLR

[AS21] Scott Armstrong and Sylvia Serfaty. Local laws and rigidity for Coulomb gases at any temperature. Annals of Probability, 49 (1):46–121, January 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-1/
Local-laws-and-rigidity-for-Coulomb-gases-at-any-temperature/
10.1214/20-AOP1445.full.

Augeri:2020:NLD

[Aug20] Fanny Augeri. Nonlinear large deviation bounds with applications to Wigner matrices and sparse Erdős–Rényi graphs. Annals of Probability, 48(5):2404–2448, September 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-5/Nonlinear-large-deviation-
bounds-with-applications-to-Wigner-matrices-and/10.1214/
20-AOP1427.full.

Banerjee:2021:RCE

[BB21a] Sayan Banerjee and Krzysztof Burdzy. Rates of convergence to equilibrium for potlatch and smoothing processes. Annals of Probability, 49(3):1129–1163, May 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-3/Rates-of-convergence-
to-equilibrium-for-potlatch-and-smoothing-processes/10.
1214/20-AOP1473.full.

Borodin:2021:CPS

[BB21b] Alexei Borodin and Alexey Bufetov. Color-position symmetry in interacting particle systems. Annals of Probability, 49(4): 1607–1632, July 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-4/
Color-position-symmetry-in-interacting-particle-systems/
10.1214/20-AOP1463.full.

Banerjee:2022:DAI

[BB22] Sayan Banerjee and Amarjit Budhiraja. Domains of attraction of invariant distributions of the infinite Atlas model. Annals of Probability, 50(4):1610–1646, July 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-4/Domains-of-attraction-
of-invariant-distributions-of-the-infinite-Atlas/10.1214/
22-AOP1570.full.

Bhattacharya:2021:SES

[BBG21] Bhaswar B. Bhattacharya, Sohom Bhattacharya, and Shirshendu Ganguly. Spectral edge in sparse random graphs: Upper and lower tail large deviations. Annals of Probability, 49(4):1847–1885, July 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-4/Spectral-edge-
in-sparse-random-graphs--Upper-and-lower/10.1214/20-AOP1495.
full.

Berestycki:2022:BBI

[BBNP22] Julien Berestycki, Éric Brunet, James Nolen, and Sarah Penington. Brownian bees in the infinite swarm limit. Annals of Probability, 50(6):2133–2177, November 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-6/Brownian-bees-in-the-
infinite-swarm-limit/10.1214/22-AOP1578.full.

Bedrossian:2022:ASE

[BBPS22] Jacob Bedrossian, Alex Blumenthal, and Samuel Punshon-Smith. Almost-sure exponential mixing of passive scalars by the stochastic Navier–Stokes equations. Annals of Probability, 50(1):241– 303, January 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-1/
Almost-sure-exponential-mixing-of-passive-scalars-by-
the-stochastic/10.1214/21-AOP1533.full.

Bates:2020:EDD

[BC20a] Erik Bates and Sourav Chatterjee. The endpoint distribution of directed polymers. Annals of Probability, 48(2):817–871, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/euclid.aop/
1587542681.

Bates:2020:LGD

[BC20b] Erik Bates and Sourav Chatterjee. Localization in Gaussian disordered systems at low temperature. Annals of Probability, 48(6):

2755–2806, November 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-6/
Localization-in-Gaussian-disordered-systems-at-low-temperature/
10.1214/20-AOP1436.full.

Bakhtin:2021:AEE

[BC21] Yuri Bakhtin and Hong-Bin Chen. Atypical exit events near a repelling equilibrium. Annals of Probability, 49(3): 1257–1285, May 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-3/
Atypical-exit-events-near-a-repelling-equilibrium/10.
1214/20-AOP1479.full.

Barraquand:2023:SML

[BC23] Guillaume Barraquand and Ivan Corwin. Stationary measures for the log-gamma polymer and KPZ equation in half-space. Annals of Probability, 51(5):1830–1869, September 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-5/Stationary-measures-
for-the-log-gamma-polymer-and-KPZ-equation/10.1214/23-
AOP1634.full.

Bailleul:2021:PCM

[BCD21] Ismaël Bailleul, Rémi Catellier, and François Delarue. Propagation of chaos for mean field rough differential equations. Annals of Probability, 49(2):944–996, March 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-2/Propagation-of-chaos-
for-mean-field-rough-differential-equations/10.1214/20-
AOP1465.full.

Berger:2022:EHI

[BCDG22] Noam Berger, Moran Cohen, Jean-Dominique Deuschel, and Xiaoqin Guo. An elliptic Harnack inequality for difference equations with random balanced coefficients. Annals of Probability, 50(3):835–873, May 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-3/
An-elliptic-Harnack-inequality-for-difference-equations-
with-random-balanced/10.1214/21-AOP1544.full.

Borgs:2021:LSC

[BCDS21] Christian Borgs, Jennifer T. Chayes, Souvik Dhara, and Subhabrata Sen. Limits of sparse configuration models and beyond: Graphexes and MultiGraphexes. Annals of Probability, 49(6): 2830–2873, November 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-6/
Limits-of-sparse-configuration-models-and-beyond--Graphexes-
and/10.1214/21-AOP1508.full.

Beckman:2021:CED

[BCE^{+}21] Erin Beckman, Keisha Cook, Nicole Eikmeier, Sarai Hernandez-Torres, and Matthew Junge. Chase-escape with death on trees. Annals of Probability, 49(5):2530–2547, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-5/Chase-escape-
with-death-on-trees/10.1214/21-AOP1514.full.

Bobkov:2020:NAW

[BCG20] S. G. Bobkov, G. P. Chistyakov, and F. Götze. Normal approximation for weighted sums under a second-order correlation condition. Annals of Probability, 48(3):1202–1219, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

https://projecteuclid.org/euclid.aop/1592359226.

Bjornberg:2022:SSS

[BCS22] Jakob Björnberg, Nicolas Curien, and Sigurdur Örn Stefánsson. Stable shredded spheres and causal random maps with large faces. Annals of Probability, 50(5):2056–2084, September 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-5/Stable-shredded-spheres-
and-causal-random-maps-with-large-faces/10.1214/22-AOP1579.
full.

Biswas:2023:FED

[BCS23] Ratul Biswas, Wei-Kuo Chen, and Arnab Sen. Free energy of a diluted spin glass model with quadratic Hamiltonian. Annals of Probability, 51(1):359–395, January 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-probability/volume-51/issue-1/Free-energy-of-a-diluted-
spin-glass-model-with-quadratic/10.1214/22-AOP1597.full.

Brito:2023:ABI

[BDH23] Gerandy Brito, Michael Damron, and Jack Hanson. Absence of backward infinite paths for first-passage percolation in arbitrary dimension. Annals of Probability, 51(1): 70–100, January 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-1/
Absence-of-backward-infinite-paths-for-first-passage-
percolation-in/10.1214/22-AOP1588.full.

Benoist:2020:NCS

[BDW20] Stéphane Benoist, Laure Dumaz, and Wendelin Werner. Nearcritical spanning forests and renormalization. Annals of Probability, 48(4):1980–2013, July 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/journals/annals-of-probability/volume-
48/issue-4/Near-critical-spanning-forests-and-renormalization/
10.1214/19-AOP1413.full.

Bender:2020:IFG

[Ben20] Christian Bender. Itô’s formula for Gaussian processes with stochastic discontinuities. Annals of Probability, 48(1):458–492, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL https://projecteuclid.org/euclid.
aop/1585123335.

Berggren:2021:DTA

[Ber21] Tomas Berggren. Domino tilings of the Aztec diamond with doubly periodic weightings. Annals of Probability, 49(4): 1965–2011, July 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-4/
Domino-tilings-of-the-Aztec-diamond-with-doubly-periodic-
weightings/10.1214/20-AOP1498.full.

Busani:2022:UGT

[BF22] Ofer Busani and Patrik L. Ferrari. Universality of the geodesic tree in last passage percolation. Annals of Probability, 50(1): 90–130, January 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-1/
Universality-of-the-geodesic-tree-in-last-passage-percolation/
10.1214/21-AOP1530.full.

Basu:2021:FGA

[BGH21] Riddhipratim Basu, Shirshendu Ganguly, and Alan Hammond. Fractal geometry of Airy. Annals of Probability, 49(1):485– 505, January 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-1/
Fractal-geometry-of-Airy_2-processes-coupled-via-the-
Airy-sheet/10.1214/20-AOP1444.full.

BenArous:2020:ATT

[BGJ20] Gérard Ben Arous, Reza Gheissari, and Aukosh Jagannath. Algorithmic thresholds for tensor PCA. Annals of Probability, 48 (4):2052–2087, July 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-4/
Algorithmic-thresholds-for-tensor-PCA/10.1214/19-AOP1415.
full.

Bertini:2023:CDV

[BGL23] Lorenzo Bertini, Davide Gabrielli, and Claudio Landim. Concurrent Donsker–Varadhan and hydrodynamical large deviations. Annals of Probability, 51(4):1298–1341, July 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-4/Concurrent-DonskerVaradhan-
and-hydrodynamical-large-deviations/10.1214/22-AOP1619.
full.

Bruno:2022:ORT

[BGW22] Stefano Bruno, Benjamin Gess, and Hendrik Weber. Optimal regularity in time and space for stochastic porous medium equations. Annals of Probability, 50(6):2288–2343, November 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-6/Optimal-regularity-in-
time-and-space-for-stochastic-porous-medium/10.1214/22-
AOP1583.full.

Bauerschmidt:2022:MCS

[BH22] Roland Bauerschmidt and Michael Hofstetter. Maximum and coupling of the sine-Gordon field. Annals of Probability, 50 (2):455–508, March 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-2/
Maximum-and-coupling-of-the-sine-Gordon-field/10.1214/
21-AOP1537.full.

Blondel:2020:RWD

[BHT20] Oriane Blondel, Marcelo R. Hilário, and Augusto Teixeira. Random walks on dynamical random environments with nonuniform mixing. Annals of Probability, 48(4):2014–2051, July 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-48/issue-4/Random-walks-
on-dynamical-random-environments-with-nonuniform-mixing/
10.1214/19-AOP1414.full.

Buraczewski:2021:DMB

[BIM21] Dariusz Buraczewski, Alexander Iksanov, and Bastien Mallein. On the derivative martingale in a branching random walk. Annals of Probability, 49(3):1164–1204, May 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-3/On-the-derivative-martingale-
in-a-branching-random-walk/10.1214/20-AOP1474.full.

Budzinski:2022:RUI

[BL22a] Thomas Budzinski and Thomas Lehéricy. Recurrence of the uniform infinite half-plane map via duality of resistances. Annals of Probability, 50(5):1725–1754, September 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-50/issue-5/Recurrence-of-
the-uniform-infinite-half-plane-map-via-duality/10.1214/
21-AOP1539.full.

Budzinski:2022:LLB

[BL22b] Thomas Budzinski and Baptiste Louf. Local limits of bipartite maps with prescribed face degrees in high genus. Annals of Probability, 50(3):1059–1126, May 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-3/Local-limits-of-bipartite-
maps-with-prescribed-face-degrees-in/10.1214/21-AOP1554.
full.

Blanquicett:2020:ABP

[Bla20] Daniel Blanquicett. Anisotropic bootstrap percolation in three dimensions. Annals of Probability, 48(5):2591–2614, September 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-48/issue-5/Anisotropic-bootstrap-
percolation-in-three-dimensions/10.1214/20-AOP1434.full.

Berestycki:2020:DIG

[BLR20] Nathanaël Berestycki, Benoît Laslier, and Gourab Ray. Dimers and imaginary geometry. Annals of Probability, 48(1):1–52, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL https://projecteuclid.org/euclid.
aop/1585123322.

Bednorz:2022:SID

[BM22a] Witold Bednorz and Rafał Martynek. The suprema of infinitely divisible processes. Annals of Probability, 50(1):397– 417, January 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-1/
The-suprema-of-infinitely-divisible-processes/10.1214/
21-AOP1546.full.

Borga:2022:SLL

[BM22b] Jacopo Borga and Mickaël Maazoun. Scaling and local limits of Baxter permutations and bipolar orientations through coalescent-walk processes. Annals of Probability, 50(4):1359– 1417, July 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-4/
Scaling-and-local-limits-of-Baxter-permutations-and-bipolar-
orientations/10.1214/21-AOP1559.full.

Beliaev:2020:CFT

[BMR20] Dmitry Beliaev, Stephen Muirhead, and Alejandro Rivera. A covariance formula for topological events of smooth Gaussian fields. Annals of Probability, 48(6):2845–2893, November 2020.

CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-6/A-covariance-formula-
for-topological-events-of-smooth-Gaussian-fields/10.1214/
20-AOP1438.full.

Bahadoran:2020:HCR

[BMRS20] C. Bahadoran, T. Mountford, K. Ravishankar, and E. Saada. Hydrodynamics in a condensation regime: The disordered asymmetric zero-range process. Annals of Probability, 48(1):404–444, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL https://projecteuclid.org/euclid.
aop/1585123333.

Banderier:2020:PPU

[BMW20] Cyril Banderier, Philippe Marchal, and Michael Wallner. Periodic Pólya urns, the density method and asymptotics of Young tableaux. Annals of Probability, 48(4):1921–1965, July 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-4/Periodic-P%c3%b3lya-
urns-the-density-method-and-asymptotics-of-Young/10.1214/
19-AOP1411.full.

Bhamidi:2021:SEE

[BNNS21] Shankar Bhamidi, Danny Nam, Oanh Nguyen, and Allan Sly. Survival and extinction of epidemics on random graphs with general degree. Annals of Probability, 49(1):244–286, January 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-1/Survival-and-
extinction-of-epidemics-on-random-graphs-with-general/
10.1214/20-AOP1451.full.

Bobkov:2023:DCD

[Bob23] Sergey G. Bobkov. Decay of convolved densities via Laplace transform. Annals of Probability, 51(5):1603–1615, September 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-5/Decay-of-convolved-densities-
via-Laplace-transform/10.1214/23-AOP1625.full.

Beffara:2021:UGM

[BPW21] Vincent Beffara, Eveliina Peltola, and Hao Wu. On the uniqueness of global multiple SLEs. Annals of Probability, 49(1): 400–434, January 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-1/
On-the-uniqueness-of-global-multiple-SLEs/10.1214/20-
AOP1477.full.

Barbu:2020:NFP

[BR20] Viorel Barbu and Michael Röckner. From nonlinear Fokker– Planck equations to solutions of distribution dependent SDE. Annals of Probability, 48(4):1902–1920, July 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-4/From-nonlinear-FokkerPlanck-
equations-to-solutions-of-distribution-dependent-SDE/
10.1214/19-AOP1410.full.

Bou-Rabee:2021:CRA

[BR21] Ahmed Bou-Rabee. Convergence of the random Abelian sandpile. Annals of Probability, 49(6):3168–3196, November 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-6/Convergence-of-the-random-
Abelian-sandpile/10.1214/21-AOP1528.full.

Bella:2020:QIP

[BS20] Peter Bella and Mathias Schäffner. Quenched invariance principle for random walks among random degenerate conductances. Annals of Probability, 48(1):296–316, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/euclid.aop/1585123329.

Bowditch:2022:TDC

[BS22] Adam Bowditch and Rongfeng Sun. The two-dimensional continuum random field Ising model. Annals of Probability, 50(2):419– 454, March 2022. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-50/issue-2/The-
two-dimensional-continuum-random-field-Ising-model/10.
1214/21-AOP1536.full.

Benjamini:2023:GIL

[BTZ23] Itai Benjamini, Hagai Helman Tov, and Maksim Zhukovskii. Global information from local observations of the noisy voter model on a graph. Annals of Probability, 51(5):1963–1992, September 2023. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-51/issue-5/Global-
information-from-local-observations-of-the-noisy-voter-
model/10.1214/23-AOP1637.full.

Backhoff-Veraguas:2020:MBB

[BVBHK20] Julio Backhoff-Veraguas, Mathias Beiglböck, Martin Huesmann, and Sigrid Källblad. Martingale Benamou–Brenier: a probabilistic perspective. Annals of Probability, 48(5):2258– 2289, September 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-5/
Martingale-BenamouBrenier-A-probabilistic-perspective/
10.1214/20-AOP1422.full.

Belius:2020:MGL

[BW20] David Belius and Wei Wu. Maximum of the Ginzburg–Landau fields. Annals of Probability, 48(6):2647–2679, November 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-6/Maximum-of-the-GinzburgLandau-
fields/10.1214/19-AOP1416.full.

Blumenthal:2023:EMR

[BZG23] Alex Blumenthal, Michele Coti Zelati, and Rishabh S. Gvalani. Exponential mixing for random dynamical systems and an example of pierrehumbert. Annals of Probability, 51(4):1559– 1601, July 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-4/
Exponential-mixing-for-random-dynamical-systems-and-an-
example-of/10.1214/23-AOP1627.full.

Chen:2021:PHN

[CCKW21] Xin Chen, Zhen-Qing Chen, Takashi Kumagai, and Jian Wang. Periodic homogenization of nonsymmetric Lévy-type processes. Annals of Probability, 49(6):2874–2921, November 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-6/Periodic-homogenization-
of-nonsymmetric-L%c3%a9vy-type-processes/10.1214/21-AOP1518.
full.

Carlen:2020:SGR

[CCL20] Eric Carlen, Maria Carvalho, and Michael Loss. Spectral gaps for reversible Markov processes with chaotic invariant measures: The Kac process with hard sphere collisions in three dimensions. Annals of Probability, 48(6):2807–2844, November 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-48/issue-6/Spectral-gaps-
for-reversible-Markov-processes-with-chaotic-invariant-
measures/10.1214/20-AOP1437.full.

Cerny:2020:QIP

[ČD20a] Jiří Černý and Alexander Drewitz. Quenched invariance principles for the maximal particle in branching random walk in random environment and the parabolic Anderson model. Annals of Probability, 48(1):94–146, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/euclid.aop/1585123324.

Chatterjee:2020:CSD

[CD20b] Sourav Chatterjee and Alexander Dunlap. Constructing a solution of the (2+1)-dimensional KPZ equation. Annals of Probability, 48 (2):1014–1055, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1587542686.

Cerf:2022:TCB

[CD22] Raphaël Cerf and Barbara Dembin. The time constant for Bernoulli percolation is Lipschitz continuous strictly above. Annals of Probability, 50(5):1781–1812, September 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-5/The-time-constant-for-
Bernoulli-percolation-is-Lipschitz-continuous-strictly/
10.1214/22-AOP1565.full.

Chen:2021:DRC

[CDD^{+}21] Xinxing Chen, Victor Dagard, Bernard Derrida, Yueyun Hu, Mikhail Lifshits, and Zhan Shi. The Derrida–Retaux conjecture on recursive models. Annals of Probability, 49(2):637– 670, March 2021. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-49/issue-2/The-
DerridaRetaux-conjecture-on-recursive-models/10.1214/
20-AOP1457.full.

Carmona:2020:EMF

[CDL20] René Carmona, François Delarue, and Daniel Lacker. Errata: Mean field games with common noise. Annals of Probability, 48(5): 2644–2646, September 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-5/
Errata-Mean-field-games-with-common-noise/10.1214/20-
AOP1432.full.

Cuny:2023:BET

[CDMP23] C. Cuny, J. Dedecker, F. Merlevède, and M. Peligrad. Berry– Esseen type bounds for the left random walk on. Annals of Probability, 51(2):495–523, March 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/journals/annals-of-probability/volume-
51/issue-2/BerryEsseen-type-bounds-for-the-left-random-
walk-on-GLdR/10.1214/22-AOP1602.full.

Chen:2021:MES

[CDOT21] Xia Chen, Aurélien Deya, Cheng Ouyang, and Samy Tindel. Moment estimates for some renormalized parabolic Anderson models. Annals of Probability, 49(5):2599–2636, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-5/Moment-estimates-for-
some-renormalized-parabolic-Anderson-models/10.1214/21-
AOP1517.full.

Comets:2023:SLH

[CDS23] Francis Comets, Joseba Dalmau, and Santiago Saglietti. Scaling limit of the heavy tailed ballistic deposition model with psticking. Annals of Probability, 51(5):1870–1931, September 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-5/Scaling-limit-of-the-
heavy-tailed-ballistic-deposition-model-with/10.1214/
23-AOP1635.full.

Chatterjee:2022:PTR

[CDSZ22] Sourav Chatterjee, Persi Diaconis, Allan Sly, and Lingfu Zhang. A phase transition for repeated averages. Annals of Probability, 50(1):1–17, January 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-1/
A-phase-transition-for-repeated-averages/10.1214/21-AOP1526.
full.

Cordero-Erausquin:2023:ILC

[CER23] Dario Cordero-Erausquin and Liran Rotem. Improved logconcavity for rotationally invariant measures of symmetric convex sets. Annals of Probability, 51(3):987–1003, May 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-51/issue-3/Improved-log-
concavity-for-rotationally-invariant-measures-of-symmetric-
convex/10.1214/22-AOP1604.full.

Cannizzaro:2021:AKS

[CES21] Giuseppe Cannizzaro, Dirk Erhard, and Philipp Schönbauer. 2D anisotropic KPZ at stationarity: Scaling, tightness and nontriviality. Annals of Probability, 49(1):122–156, January 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-1/2D-anisotropic-KPZ-at-
stationarity-Scaling-tightness-and-nontriviality/10.1214/
20-AOP1446.full.

Cipolloni:2022:NFQ

[CES22] Giorgio Cipolloni, László Erdős, and Dominik Schröder. Normal fluctuation in quantum ergodicity for Wigner matrices. Annals of Probability, 50(3):984–1012, May 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-3/Normal-fluctuation-in-
quantum-ergodicity-for-Wigner-matrices/10.1214/21-AOP1552.
full.

Corwin:2021:KEC

[CGH21] Ivan Corwin, Promit Ghosal, and Alan Hammond. KPZ equation correlations in time. Annals of Probability, 49(2):832– 876, March 2021. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-49/issue-2/KPZ-
equation-correlations-in-time/10.1214/20-AOP1461.full.

Cryan:2021:MLS

[CGM21] Mary Cryan, Heng Guo, and Giorgos Mousa. Modified log-Sobolev inequalities for strongly log-concave distributions. Annals of Probability, 49(1):506–525, January 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-1/Modified-log-Sobolev-
inequalities-for-strongly-log-concave-distributions/10.
1214/20-AOP1453.full.

Chatterjee:2020:LRG

[CH20] Sourav Chatterjee and Matan Harel. Localization in random geometric graphs with too many edges. Annals of Probability, 48 (2):574–621, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1587542673.

Chelkak:2023:USC

[CIM23] Dmitry Chelkak, Konstantin Izyurov, and Rémy Mahfouf. Universality of spin correlations in the Ising model on isoradial graphs. Annals of Probability, 51(3):840–898, May 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-3/Universality-of-spin-
correlations-in-the-Ising-model-on-isoradial/10.1214/
22-AOP1595.full.

Chong:2020:ASA

[CK20] Carsten Chong and Péter Kevei. The almost-sure asymptotic behavior of the solution to the stochastic heat equation with Lévy noise. Annals of Probability, 48(3):1466–1494, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

https://projecteuclid.org/euclid.aop/1592359235.

Conchon-Kerjan:2022:CRL

[CK22] Guillaume Conchon-Kerjan. Cutoff for random lifts of weighted graphs. Annals of Probability, 50(1):304–338, January 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-probability/volume-50/issue-1/Cutoff-for-random-lifts-
of-weighted-graphs/10.1214/21-AOP1534.full.

Chong:2023:LPI

[CK23] Carsten Chong and Péter Kevei. A landscape of peaks: The intermittency islands of the stochastic heat equation with Lévy noise. Annals of Probability, 51(4):1449–1501, July 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-4/A-landscape-of-peaks-
-The-intermittency-islands-of-the/10.1214/23-AOP1623.
full.

Conchon-Kerjan:2023:SGM

[CKG23] Guillaume Conchon-Kerjan and Christina Goldschmidt. The stable graph: The metric space scaling limit of a critical random graph with i.i.d. power-law degrees. Annals of Probability, 51(1): 1–69, January 2023. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-51/issue-1/The-
stable-graph--The-metric-space-scaling-limit-of/10.1214/
22-AOP1587.full.

Caputo:2020:MTA

[CLL20] Pietro Caputo, Cyril Labbé, and Hubert Lacoin. Mixing time of the adjacent walk on the simplex. Annals of Probability, 48(5): 2449–2493, September 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-5/
Mixing-time-of-the-adjacent-walk-on-the-simplex/10.1214/
20-AOP1428.full.

Chen:2023:LHK

[CLW23] Xin Chen, Xue-Mei Li, and Bo Wu. Logarithmic heat kernel estimates without curvature restrictions. Annals of Probability, 51 (2):442–477, March 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-2/
Logarithmic-heat-kernel-estimates-without-curvature-restrictions/
10.1214/22-AOP1599.full.

Chiarini:2020:ERO

[CN20] Alberto Chiarini and Maximilian Nitzschner. Entropic repulsion for the occupation-time field of random interlacements conditioned on disconnection. Annals of Probability, 48(3):1317–1351, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL https://projecteuclid.org/euclid.
aop/1592359230.

Campese:2020:CBM

[CNN20] Simon Campese, Ivan Nourdin, and David Nualart. Continuous Breuer–Major theorem: Tightness and nonstationarity. Annals of Probability, 48(1):147–177, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/euclid.aop/1585123325.

Chasapis:2022:S

[CNT22] Giorgos Chasapis, Piotr Nayar, and Tomasz Tkocz. Slicing. Annals of Probability, 50(6):2344–2372, November 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-6/Slicing-%e2%84%93p-balls-
reloaded-Stability-planar-sections-in-%e2%84%931/10.1214/
22-AOP1584.full.

Crane:2021:AEM

[CRY21] Edward Crane, Balázs Ráth, and Dominic Yeo. Age evolution in the mean field forest fire model via multitype branching processes. Annals of Probability, 49(4):2031–2075, July 2021. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-49/issue-4/Age-
evolution-in-the-mean-field-forest-fire-model-via/10.
1214/20-AOP1501.full.

Coste:2021:EES

[CS21] Simon Coste and Justin Salez. Emergence of extended states at zero in the spectrum of sparse random graphs. Annals of Probability, 49(4):2012–2030, July 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-4/Emergence-of-extended-
states-at-zero-in-the-spectrum-of/10.1214/20-AOP1499.
full.

Coupier:2021:DSF

[CSST21] David Coupier, Kumarjit Saha, Anish Sarkar, and Viet Chi Tran. The 2d-directed spanning forest converges to the Brownian web. Annals of Probability, 49(1):435–484, January 2021. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-49/issue-1/The-
2d-directed-spanning-forest-converges-to-the-Brownian-
web/10.1214/20-AOP1478.full.

Caravenna:2020:TDK

[CSZ20] Francesco Caravenna, Rongfeng Sun, and Nikos Zygouras. The two-dimensional KPZ equation in the entire subcritical regime. Annals of Probability, 48(3):1086–1127, May 2020. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

https://projecteuclid.org/euclid.aop/1592359223.

Chouk:2021:AEA

[CvZ21] Khalil Chouk and Willem van Zuijlen. Asymptotics of the eigenvalues of the Anderson Hamiltonian with white noise potential in two dimensions. Annals of Probability, 49(4): 1917–1964, July 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-4/
Asymptotics-of-the-eigenvalues-of-the-Anderson-Hamiltonian-
with-white/10.1214/20-AOP1497.full.

Cerrai:2022:SKA

[CX22] Sandra Cerrai and Guangyu Xi. A Smoluchowski–Kramers approximation for an infinite dimensional system with statedependent damping. Annals of Probability, 50(3):874–904, May 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-50/issue-3/A-SmoluchowskiKramers-
approximation-for-an-infinite-dimensional-system-with-
state/10.1214/21-AOP1549.full.

Dauvergne:2022:HIL

[Dau22] Duncan Dauvergne. Hidden invariance of last passage percolation and directed polymers. Annals of Probability, 50(1): 18–60, January 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-1/
Hidden-invariance-of-last-passage-percolation-and-directed-
polymers/10.1214/21-AOP1527.full.

Duminil-Copin:2020:BNS

[DCGHM20] Hugo Duminil-Copin, Shirshendu Ganguly, Alan Hammond, and Ioan Manolescu. Bounding the number of self-avoiding walks: Hammersley–Welsh with polygon insertion. Annals of Probability, 48(4):1644–1692, July 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-4/
Bounding-the-number-of-self-avoiding-walks--HammersleyWelsh-
with/10.1214/19-AOP1400.full.

Duminil-Copin:2020:NMP

[DCKN^{+}20] Hugo Duminil-Copin, Harry Kesten, Fedor Nazarov, Yuval Peres, and Vladas Sidoravicius. On the number of maximal paths in directed last-passage percolation. Annals of Probability, 48(5): 2176–2188, September 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-5/
On-the-number-of-maximal-paths-in-directed-last-passage/
10.1214/19-AOP1419.full.

Duminil-Copin:2023:EUN

[DCRRV23] Hugo Duminil-Copin, Alejandro Rivera, Pierre-François Rodriguez, and Hugo Vanneuville. Existence of an unbounded nodal hypersurface for smooth Gaussian fields in dimension. Annals of Probability, 51(1):228–276, January 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-1/Existence-of-an-unbounded-
nodal-hypersurface-for-smooth-Gaussian-fields/10.1214/
22-AOP1594.full.

Dembin:2020:MFC

[Dem20] Barbara Dembin. The maximal flow from a compact convex subset to infinity in first passage percolation on Z^{d}. Annals of Probability, 48(2):622–645, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1587542674.

Ding:2021:DRW

[DFSX21] Jian Ding, Ryoki Fukushima, Rongfeng Sun, and Changji Xu. Distribution of the random walk conditioned on survival among quenched Bernoulli obstacles. Annals of Probability, 49(1): 206–243, January 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-1/
Distribution-of-the-random-walk-conditioned-on-survival-
among-quenched/10.1214/20-AOP1450.full.

Dario:2021:QHP

[DG21] Paul Dario and Chenlin Gu. Quantitative homogenization of the parabolic and elliptic Green’s functions on percolation clusters. Annals of Probability, 49(2):556–636, March 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-2/Quantitative-
homogenization-of-the-parabolic-and-elliptic-Greens-functions-
on/10.1214/20-AOP1456.full.

Dunlap:2022:FBS

[DG22] Alexander Dunlap and Yu Gu. A forward-backward SDE from the 2D nonlinear stochastic heat equation. Annals of Probability, 50(3):1204–1253, May 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-3/
A-forward-backward-SDE-from-the-2D-nonlinear-stochastic-
heat/10.1214/21-AOP1563.full.

Das:2023:LIL

[DG23] Sayan Das and Promit Ghosal. Law of iterated logarithms and fractal properties of the KPZ equation. Annals of Probability, 51(3):930–986, May 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-3/
Law-of-iterated-logarithms-and-fractal-properties-of-
the-KPZ/10.1214/22-AOP1603.full.

Dyszewski:2022:SCL

[DGJ^{+}22] Piotr Dyszewski, Nina Gantert, Samuel G. G. Johnston, Joscha Prochno, and Dominik Schmid. Sharp concentration for the largest and smallest fragment in a k-regular self-similar fragmentation. Annals of Probability, 50(3):1173–1203, May 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-3/Sharp-concentration-
for-the-largest-and-smallest-fragment-in-a/10.1214/21-
AOP1556.full.

Diaconu:2023:MLD

[Dia23] Simona Diaconu. More limiting distributions for eigenvalues of Wigner matrices. Annals of Probability, 51(2):774– 804, March 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-2/
More-limiting-distributions-for-eigenvalues-of-Wigner-
matrices/10.1214/22-AOP1614.full.

Doring:2020:EEI

[DK20] Leif Döring and Andreas E. Kyprianou. Entrance and exit at infinity for stable jump diffusions. Annals of Probability, 48 (3):1220–1265, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1592359227.

Devraj:2020:GEW

[DKM20] Adithya Devraj, Ioannis Kontoyiannis, and Sean Meyn. Geometric ergodicity in a weighted Sobolev space. Annals of Probability, 48 (1):380–403, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1585123332.

Dumaz:2023:DPA

[DL23] Laure Dumaz and Cyril Labbé. The delocalized phase of the Anderson Hamiltonian in 1-D. Annals of Probability, 51(3):805– 839, May 2023. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-51/issue-3/The-
delocalized-phase-of-the-Anderson-Hamiltonian-in-1-D/
10.1214/22-AOP1591.full.

Delarue:2020:MEM

[DLR20] François Delarue, Daniel Lacker, and Kavita Ramanan. From the master equation to mean field game limit theory: Large deviations and concentration of measure. Annals of Probability, 48 (1):211–263, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1585123327.

Dudeja:2023:UAM

[DLS23] Rishabh Dudeja, Yue M. Lu, and Subhabrata Sen. Universality of approximate message passing with semirandom matrices. Annals of Probability, 51(5):1616–1683, September 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-51/issue-5/Universality-
of-approximate-message-passing-with-semirandom-matrices/
10.1214/23-AOP1628.full.

Dimitrov:2021:CBG

[DM21] Evgeni Dimitrov and Konstantin Matetski. Characterization of Brownian Gibbsian line ensembles. Annals of Probability, 49(5): 2477–2529, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-5/
Characterization-of-Brownian-Gibbsian-line-ensembles/
10.1214/21-AOP1513.full.

Dalang:2021:PAA

[DMX21] Robert C. Dalang, Carl Mueller, and Yimin Xiao. Polarity of almost all points for systems of nonlinear stochastic heat equations in the critical dimension. Annals of Probability, 49(5): 2573–2598, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-5/
Polarity-of-almost-all-points-for-systems-of-nonlinear-
stochastic/10.1214/21-AOP1516.full.

Deuschel:2021:AFR

[DOP21] Jean-Dominique Deuschel, Tal Orenshtein, and Nicolas Perkowski. Additive functionals as rough paths. Annals of Probability, 49 (3):1450–1479, May 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-3/
Additive-functionals-as-rough-paths/10.1214/20-AOP1488.
full.

Dubach:2021:WNH

[DP21] Guillaume Dubach and Yuval Peled. On words of non-Hermitian random matrices. Annals of Probability, 49(4): 1886–1916, July 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-4/
On-words-of-non-Hermitian-random-matrices/10.1214/20-
AOP1496.full.

Djete:2022:MVO

[DPT22] Mao Fabrice Djete, Dylan Possamaï, and Xiaolu Tan. McKean– Vlasov optimal control: The dynamic programming principle. Annals of Probability, 50(2):791–833, March 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-2/McKeanVlasov-optimal-
control-The-dynamic-programming-principle/10.1214/21-
AOP1548.full.

deRaphelis:2022:SLS

[dR22] Loïc de Raphélis. Scaling limit of the subdiffusive random walk on a Galton–Watson tree in random environment. Annals of Probability, 50(1):339–396, January 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-1/Scaling-limit-of-the-
subdiffusive-random-walk-on-a-GaltonWatson/10.1214/21-
AOP1535.full.

Dauvergne:2022:THV

[DSV22] Duncan Dauvergne, Sourav Sarkar, and Bálint Virág. Threehalves variation of geodesics in the directed landscape. Annals of Probability, 50(5):1947–1985, September 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-5/Three-halves-variation-
of-geodesics-in-the-directed-landscape/10.1214/22-AOP1574.
full.

Dey:2023:SMC

[DT23] Partha S. Dey and Grigory Terlov. Stein’s method for conditional central limit theorem. Annals of Probability, 51(2):723–773, March 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-51/issue-2/Steins-method-
for-conditional-central-limit-theorem/10.1214/22-AOP1613.
full.

Dette:2020:RMP

[DTV20] Holger Dette, Dominik Tomecki, and Martin Venker. Random moment problems under constraints. Annals of Probability, 48 (2):672–713, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1587542676.

Dauvergne:2021:BPA

[DV21] Duncan Dauvergne and Bálint Virág. Bulk properties of the Airy line ensemble. Annals of Probability, 49(4):1738– 1777, July 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-4/
Bulk-properties-of-the-Airy-line-ensemble/10.1214/20-
AOP1492.full.

Dereudre:2023:NRB

[DV23] David Dereudre and Thibaut Vasseur. Number-rigidity and β-circular Riesz gas. Annals of Probability, 51(3):1025– 1065, May 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-3/
Number-rigidity-and-%ce%b2-circular-Riesz-gas/10.1214/
22-AOP1606.full.

Ding:2020:PLS

[DW20] Jian Ding and Mateo Wirth. Percolation for level-sets of Gaussian free fields on metric graphs. Annals of Probability, 48(3): 1411–1435, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1592359233.

Eisenbaum:2014:CPC

[Eis14] Nathalie Eisenbaum. Characterization of positively correlated squared Gaussian processes. Annals of Probability, 42(2):559–575, March 2014. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL http://projecteuclid.org/euclid.

aop/1393251296. See erratum [Eis22].

Eisenbaum:2022:ECP

[Eis22] Nathalie Eisenbaum. Erratum to “Characterization of positively correlated squared Gaussian processes”. Annals of Probability, 50(4):1674–??, July 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-4/
Erratum-to-Characterization-of-positively-correlated-
squared-Gaussian-processes/10.1214/21-AOP1561.full. See

[Eis14].

Erde:2023:ESR

[EKK23] Joshua Erde, Mihyun Kang, and Michael Krivelevich. Expansion in supercritical random subgraphs of the hypercube and its consequences. Annals of Probability, 51(1):127–156, January 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-51/issue-1/Expansion-in-
supercritical-random-subgraphs-of-the-hypercube-and-its/
10.1214/22-AOP1592.full.

ElAlaoui:2021:OMF

[EMS21] Ahmed El Alaoui, Andrea Montanari, and Mark Sellke. Optimization of mean-field spin glasses. Annals of Probability, 49(6): 2922–2960, November 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-6/
Optimization-of-mean-field-spin-glasses/10.1214/21-AOP1519.
full.

Eldan:2020:CHD

[EMZ20] Ronen Eldan, Dan Mikulincer, and Alex Zhai. The CLT in high dimensions: Quantitative bounds via martingale embedding. Annals of Probability, 48(5):2494–2524, September 2020. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-48/issue-5/The-
CLT-in-high-dimensions--Quantitative-bounds-via-martingale/
10.1214/20-AOP1429.full.

Feldheim:2021:PGS

[FFN21] Naomi Feldheim, Ohad Feldheim, and Shahaf Nitzan. Persistence of Gaussian stationary processes: A spectral perspective. Annals of Probability, 49(3):1067–1096, May 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-3/Persistence-of-Gaussian-
stationary-processes-A-spectral-perspective/10.1214/20-
AOP1470.full.

Friz:2022:FCM

[FGR22] Peter K. Friz, Jim Gatheral, and Radoš Radoičić. Forests, cumulants, martingales. Annals of Probability, 50(4):1418– 1445, July 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-4/
Forests-cumulants-martingales/10.1214/21-AOP1560.full.

Freidlin:2023:PPE

[FK23] Mark Freidlin and Leonid Koralov. Perturbations of parabolic equations and diffusion processes with degeneration: Boundary problems, metastability, and homogenization. Annals of Probability, 51(5):1752–1784, September 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/journals/annals-of-probability/volume-
51/issue-5/Perturbations-of-parabolic-equations-and-diffusion-
processes-with-degeneration/10.1214/23-AOP1631.full.

Fox:2021:ACS

[FKS21] Jacob Fox, Matthew Kwan, and Lisa Sauermann. Anticoncentration for subgraph counts in random graphs. Annals of Probability, 49(3):1515–1553, May 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/journals/annals-of-probability/volume-
49/issue-3/Anti-concentration-for-subgraph-counts-in-
random-graphs/10.1214/20-AOP1490.full.

Fox:2023:GML

[FKS23] Jacob Fox, Matthew Kwan, and Hunter Spink. Geometric and o-minimal Littlewood–Offord problems. Annals of Probability, 51 (1):101–126, January 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-1/
Geometric-and-o-minimal-LittlewoodOfford-problems/10.
1214/22-AOP1590.full.

Flandoli:2020:CTN

[FL20] Franco Flandoli and Dejun Luo. Convergence of transport noise to Ornstein–Uhlenbeck for 2D Euler equations under the enstrophy measure. Annals of Probability, 48(1):264–295, January 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL https://projecteuclid.org/euclid.aop/1585123328.

Fan:2021:TFE

[FMM21] Zhou Fan, Song Mei, and Andrea Montanari. TAP free energy, spin glasses and variational inference. Annals of Probability, 49(1): 1–45, January 2021. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-49/issue-1/TAP-
free-energy-spin-glasses-and-variational-inference/10.
1214/20-AOP1443.full.

Forman:2021:DSI

[FPRW21] Noah Forman, Soumik Pal, Douglas Rizzolo, and Matthias Winkel. Diffusions on a space of interval partitions: Poisson– Dirichlet stationary distributions. Annals of Probability, 49(2): 793–831, March 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-2/
Diffusions-on-a-space-of-interval-partitions--PoissonDirichlet-
stationary/10.1214/20-AOP1460.full.

Fournier:2020:ADM

[FT20] Nicolas Fournier and Camille Tardif. Anomalous diffusion for multi-dimensional critical kinetic Fokker–Planck equations. Annals of Probability, 48(5):2359–2403, September 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-5/Anomalous-diffusion-
for-multi-dimensional-critical-kinetic-FokkerPlanck-equations/
10.1214/20-AOP1426.full.

Forghani:2022:STL

[FT22] Behrang Forghani and Giulio Tiozzo. Shannon’s theorem for locally compact groups. Annals of Probability, 50(1): 61–89, January 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-1/
Shannons-theorem-for-locally-compact-groups/10.1214/21-
AOP1529.full.

FitzGerald:2022:AEC

[FTZ22] Will FitzGerald, Roger Tribe, and Oleg Zaboronski. Asymptotic expansions for a class of Fredholm Pfaffians and interacting particle systems. Annals of Probability, 50(6):2409– 2474, November 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-6/
Asymptotic-expansions-for-a-class-of-Fredholm-Pfaffians-
and-interacting/10.1214/22-AOP1586.full.

Feng:2021:SGC

[FW21] Renjie Feng and Dongyi Wei. Small gaps of circular βensemble. Annals of Probability, 49(2):997–1032, March 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-2/Small-gaps-of-
circular-%ce%b2-ensemble/10.1214/20-AOP1468.full.

Friz:2023:RSV

[FZK23] Peter K. Friz and Pavel Zorin-Kranich. Rough semimartingales and p-variation estimates for martingale transforms. Annals of Probability, 51(2):397–441, March 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-2/Rough-semimartingales-
and-p-variation-estimates-for-martingale-transforms/10.
1214/22-AOP1598.full.

Galashin:2021:SSC

[Gal21] Pavel Galashin. Symmetries of stochastic colored vertex models. Annals of Probability, 49(5):2175–2219, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-5/Symmetries-of-
stochastic-colored-vertex-models/10.1214/20-AOP1502.full.

Gerstenberg:2020:EIH

[Ger20] Julian Gerstenberg. Exchangeable interval hypergraphs and limits of ordered discrete structures. Annals of Probability, 48(3): 1128–1167, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1592359224.

Ganguly:2021:LGG

[GG21] Shirshendu Ganguly and Reza Gheissari. Local and global geometry of the 2D Ising interface in critical prewetting. Annals of Probability, 49(4):2076–2140, July 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-4/Local-and-global-geometry-
of-the-2D-Ising-interface-in/10.1214/21-AOP1505.full.

Giunti:2022:QHI

[GGM22] Arianna Giunti, Chenlin Gu, and Jean-Christophe Mourrat. Quantitative homogenization of interacting particle systems. Annals of Probability, 50(5):1885–1946, September 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-50/issue-5/Quantitative-
homogenization-of-interacting-particle-systems/10.1214/
22-AOP1573.full.

Guionnet:2020:LDL

[GH20] Alice Guionnet and Jonathan Husson. Large deviations for the largest eigenvalue of Rademacher matrices. Annals of Probability, 48(3):1436–1465, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1592359234.

Goldman:2022:FRD

[GH22] Michael Goldman and Martin Huesmann. A fluctuation result for the displacement in the optimal matching problem. Annals of Probability, 50(4):1446–1477, July 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-4/A-fluctuation-result-
for-the-displacement-in-the-optimal-matching/10.1214/
21-AOP1562.full.

Gwynne:2020:ASK

[GHM20] Ewain Gwynne, Nina Holden, and Jason Miller. An almost sure KPZ relation for SLE and Brownian motion. Annals of Probability, 48(2):527–573, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1587542672.

Gamarnik:2021:OGP

[GJ21] David Gamarnik and Aukosh Jagannath. The overlap gap property and approximate message passing algorithms for. Annals of Probability, 49(1):180–205, January 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-1/The-overlap-gap-property-
and-approximate-message-passing-algorithms-for/10.1214/
20-AOP1448.full.

Ghoussoub:2021:SMT

[GKP21] Nassif Ghoussoub, Young-Heon Kim, and Aaron Zeff Palmer. A solution to the Monge transport problem for Brownian martingales. Annals of Probability, 49(2):877–907, March 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-2/A-solution-to-
the-Monge-transport-problem-for-Brownian-martingales/
10.1214/20-AOP1462.full.

Gheissari:2021:TTM

[GL21] Reza Gheissari and Eyal Lubetzky. Tightness and tails of the maximum in 3D Ising interfaces. Annals of Probability, 49(2): 732–792, March 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-2/
Tightness-and-tails-of-the-maximum-in-3D-Ising-interfaces/
10.1214/20-AOP1459.full.

Giacomin:2022:DLF

[GL22]  Giambattista Giacomin and Hubert Lacoin.  The disor-

dered lattice free field pinning model approaching criticality. Annals of Probability, 50(4):1478–1537, July 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-4/The-disordered-lattice-
free-field-pinning-model-approaching-criticality/10.1214/
22-AOP1566.full.

Gwynne:2020:CGL

[GM20] Ewain Gwynne and Jason Miller. Confluence of geodesics in Liouville quantum gravity for. Annals of Probability, 48(4): 1861–1901, July 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-4/
Confluence-of-geodesics-in-Liouville-quantum-gravity-
for-gamma-in/10.1214/19-AOP1409.full.

Gwynne:2021:RWR

[GM21] Ewain Gwynne and Jason Miller. Random walk on random planar maps: Spectral dimension, resistance and displacement. Annals of Probability, 49(3):1097–1128, May 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-3/Random-walk-on-random-
planar-maps--Spectral-dimension-resistance/10.1214/20-
AOP1471.full.

Gangbo:2022:MFG

[GMMZ22] Wilfrid Gangbo, Alpár R. Mészáros, Chenchen Mou, and Jianfeng Zhang. Mean field games master equations with nonseparable Hamiltonians and displacement monotonicity. Annals of Probability, 50(6):2178–2217, November 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-6/Mean-field-games-master-
equations-with-nonseparable-Hamiltonians-and-displacement/
10.1214/22-AOP1580.full.

Gwynne:2021:TEM

[GMS21] Ewain Gwynne, Jason Miller, and Scott Sheffield. The Tutte embedding of the mated-CRT map converges to Liouville quantum gravity. Annals of Probability, 49(4):1677–1717, July 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-4/The-Tutte-embedding-
of-the-mated-CRT-map-converges-to/10.1214/20-AOP1487.
full.

Geng:2022:PLE

[GOT22] Xi Geng, Cheng Ouyang, and Samy Tindel. Precise local estimates for differential equations driven by fractional Brownian motion: Hypoelliptic case. Annals of Probability, 50(2):649–687, March 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-50/issue-2/Precise-local-
estimates-for-differential-equations-driven-by-fractional-
Brownian/10.1214/21-AOP1542.full.

Gwynne:2020:CPA

[GP20] Ewain Gwynne and Joshua Pfeffer. Connectivity properties of the adjacency graph of SLE_{κ} bubbles for κ ∈ (4, 8). Annals of Probability, 48(3):1495–1519, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/euclid.aop/1592359236.

Gwynne:2021:EDL

[GP21] Ewain Gwynne and Joshua Pfeffer. External diffusion-limited aggregation on a spanning-tree-weighted random planar map. Annals of Probability, 49(4):1633–1676, July 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-4/External-diffusion-limited-
aggregation-on-a-spanning-tree-weighted-random/10.1214/
20-AOP1486.full.

Goswami:2022:RGF

[GRS22] Subhajit Goswami, Pierre-François Rodriguez, and Franco Severo. On the radius of Gaussian free field excursion clusters. Annals of Probability, 50(5):1675–1724, September 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-50/issue-5/On-the-radius-
of-Gaussian-free-field-excursion-clusters/10.1214/22-
AOP1569.full.

Halberstam:2023:MTR

[HH23] Noah Halberstam and Tom Hutchcroft. Most transient random walks have infinitely many cut times. Annals of Probability, 51(5): 1932–1962, September 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-5/
Most-transient-random-walks-have-infinitely-many-cut-
times/10.1214/23-AOP1636.full.

Harris:2022:YLC

[HHKW22] Simon C. Harris, Emma Horton, Andreas E. Kyprianou, and Minmin Wang. Yaglom limit for critical nonlocal branching Markov processes. Annals of Probability, 50(6):2373–2408, November 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-50/issue-6/Yaglom-limit-
for-critical-nonlocal-branching-Markov-processes/10.1214/
22-AOP1585.full.

Holroyd:2020:MPF

[HHL20] Alexander E. Holroyd, Tom Hutchcroft, and Avi Levy. Mallows permutations and finite dependence. Annals of Probability, 48 (1):343–379, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1585123331.

Hutzenthaler:2020:PTS

[HJ20] Martin Hutzenthaler and Arnulf Jentzen. On a perturbation theory and on strong convergence rates for stochastic ordinary and partial differential equations with nonglobally monotone coefficients. Annals of Probability, 48(1):53–93, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

https://projecteuclid.org/euclid.aop/1585123323.

Hairer:2020:ADD

[HL20] Martin Hairer and Xue-Mei Li. Averaging dynamics driven by fractional Brownian motion. Annals of Probability, 48(4): 1826–1860, July 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-4/
Averaging-dynamics-driven-by-fractional-Brownian-motion/
10.1214/19-AOP1408.full.

Huang:2020:TTW

[HLY20] Jiaoyang Huang, Benjamin Landon, and Horng-Tzer Yau. Transition from Tracy–Widom to Gaussian fluctuations of extremal eigenvalues of sparse Erdős–Rényi graphs. Annals of Probability, 48(2):916–962, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1587542683.

Hermon:2022:MFB

[HLYZ22] Jonathan Hermon, Shuangping Li, Dong Yao, and Lingfu Zhang. Mean field behavior during the Big Bang regime for coalescing random walks. Annals of Probability, 50(5):1813–1884, September 2022. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-50/issue-5/Mean-
field-behavior-during-the-Big-Bang-regime-for-coalescing/
10.1214/22-AOP1571.full.

Hong:2020:TBR

[HMP20] Jieliang Hong, Leonid Mytnik, and Edwin Perkins. On the topological boundary of the range of super-Brownian motion. Annals of Probability, 48(3):1168–1201, May 2020. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/euclid.aop/1592359225.

Hutchcroft:2023:HDN

[HMS23] Tom Hutchcroft, Emmanuel Michta, and Gordon Slade. Highdimensional near-critical percolation and the torus plateau. Annals of Probability, 51(2):580–625, March 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-2/High-dimensional-near-
critical-percolation-and-the-torus-plateau/10.1214/22-
AOP1608.full.

Hartarsky:2021:UCK

[HMT21] Ivailo Hartarsky, Fabio Martinelli, and Cristina Toninelli. Universality for critical KCM: Finite number of stable directions. Annals of Probability, 49(5):2141–2174, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-5/Universality-
for-critical-KCM-Finite-number-of-stable-directions/10.
1214/20-AOP1500.full.

Hermon:2020:EPM

[HP20] Jonathan Hermon and Richard Pymar. The exclusion process mixes (almost) faster than independent particles. Annals of Probability, 48(6):3077–3123, November 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-6/The-exclusion-process-
mixes-almost-faster-than-independent-particles/10.1214/
20-AOP1455.full.

Holden:2021:GAU

[HPZ21] Nina Holden, Yuval Peres, and Alex Zhai. Gravitational allocation for uniform points on the sphere. Annals of Probability, 49 (1):287–321, January 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-1/
Gravitational-allocation-for-uniform-points-on-the-sphere/
10.1214/20-AOP1452.full.

Hermon:2020:CMF

[HS20a] Jonathan Hermon and Justin Salez. Cutoff for the mean-field zero-range process with bounded monotone rates. Annals of Probability, 48(2):742–759, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https:/

/projecteuclid.org/euclid.aop/1587542678.

Hermon:2020:CPR

[HS20b] Jonathan Hermon and Perla Sousi. A comparison principle for random walk on dynamical percolation. Annals of Probability, 48(6): 2952–2987, November 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-6/
A-comparison-principle-for-random-walk-on-dynamical-percolation/
10.1214/20-AOP1441.full.

Holmes:2021:STO

[HS21a] Mark Holmes and Thomas S. Salisbury. A shape theorem for the orthant model. Annals of Probability, 49(3):1237–1256, May 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-3/A-shape-theorem-for-
the-orthant-model/10.1214/20-AOP1476.full.

Hough:2021:CST

[HS21b] Robert Hough and Hyojeong Son. Cut-off for sandpiles on tiling graphs. Annals of Probability, 49(2):671–731, March 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-2/Cut-off-for-sandpiles-
on-tiling-graphs/10.1214/20-AOP1458.full.

Hammersley:2021:WEU

[HŠS21] William R. P. Hammersley, David Šiška, and Łukasz Szpruch. Weak existence and uniqueness for McKean–Vlasov SDEs with common noise. Annals of Probability, 49(2):527–555, March 2021. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-49/issue-2/Weak-
existence-and-uniqueness-for-McKeanVlasov-SDEs-with-common-
noise/10.1214/20-AOP1454.full.

Hermon:2022:UCG

[HSS22] Jonathan Hermon, Allan Sly, and Perla Sousi. Universality of cutoff for graphs with an added random matching. Annals of Probability, 50(1):203–240, January 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-1/Universality-of-cutoff-
for-graphs-with-an-added-random-matching/10.1214/21-AOP1532.
full.

Hutchcroft:2020:LCP

[Hut20] Tom Hutchcroft. Locality of the critical probability for transitive graphs of exponential growth. Annals of Probability, 48 (3):1352–1371, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1592359231.

Hofmanova:2023:GTP

[HZZ23] Martina Hofmanová, Rongchan Zhu, and Xiangchan Zhu. Globalin-time probabilistically strong and Markov solutions to stochastic 3D Navier–Stokes equations: Existence and nonuniqueness. Annals of Probability, 51(2):524–579, March 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-2/Global-in-time-probabilistically-
strong-and-Markov-solutions-to-stochastic/10.1214/22-
AOP1607.full.

Ioffe:2022:CPI

[IOSV22] Dmitry Ioffe, Sébastien Ott, Senya Shlosman, and Yvan Velenik. Critical prewetting in the 2d Ising model. Annals of Probability, 50(3):1127–1172, May 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-3/
Critical-prewetting-in-the-2d-Ising-model/10.1214/21-
AOP1555.full.

Izyurov:2022:MSF

[Izy22] Konstantin Izyurov. On multiple SLE for the FK–Ising model. Annals of Probability, 50(2):771–790, March 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-probability/volume-50/issue-2/On-multiple-SLE-for-
the-FKIsing-model/10.1214/21-AOP1547.full.

Jego:2020:PBM

[Jeg20] Antoine Jego. Planar Brownian motion and Gaussian multiplicative chaos. Annals of Probability, 48(4):1597–1643, July 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-4/Planar-Brownian-motion-
and-Gaussian-multiplicative-chaos/10.1214/19-AOP1399.
full.

Johansson:2021:MNA

[JL21] Kurt Johansson and Gaultier Lambert. Multivariate normal approximation for traces of random unitary matrices. Annals of Probability, 49(6):2961–3010, November 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-6/Multivariate-normal-
approximation-for-traces-of-random-unitary-matrices/10.
1214/21-AOP1520.full.

Janson:2023:UMV

[JL23] Svante Janson and Baptiste Louf. Unicellular maps vs. hyperbolic surfaces in large genus: Simple closed curves. Annals of Probability, 51(3):899–929, May 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-3/
Unicellular-maps-vs-hyperbolic-surfaces-in-large-genus-
-Simple/10.1214/22-AOP1601.full.

Jara:2020:SER

[JM20] Milton Jara and Otávio Menezes. Symmetric exclusion as a random environment: Invariance principle. Annals of Probability, 48(6):3124–3149, November 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/journals/annals-of-probability/volume-
48/issue-6/Symmetric-exclusion-as-a-random-environment-
Invariance-principle/10.1214/20-AOP1466.full.

Jaramillo:2023:LTA

[JNNP23] Arturo Jaramillo, Ivan Nourdin, David Nualart, and Giovanni Peccati. Limit theorems for additive functionals of the fractional Brownian motion. Annals of Probability, 51(3): 1066–1111, May 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-3/
Limit-theorems-for-additive-functionals-of-the-fractional-
Brownian-motion/10.1214/22-AOP1612.full.

Johansson:2022:IPG

[JR22]  Kurt Johansson and Mustazee Rahman.  On inhomoge-

neous polynuclear growth. Annals of Probability, 50(2):559– 590, March 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-2/
On-inhomogeneous-polynuclear-growth/10.1214/21-AOP1540.
full.

Janjigian:2020:BFG

[JRA20] Christopher Janjigian and Firas Rassoul-Agha. Busemann functions and Gibbs measures in directed polymer models on Z^{2}. Annals of Probability, 48(2):778–816, March 2020. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

https://projecteuclid.org/euclid.aop/1587542680.

Janvresse:2020:EPS

[JRdlR20] Élise Janvresse, Emmanuel Roy, and Thierry de la Rue. Ergodic Poisson splittings. Annals of Probability, 48(3):1266–1285, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL https://projecteuclid.org/euclid.
aop/1592359228.

Jain:2021:RDC

[JSS21] Vishesh Jain, Ashwin Sah, and Mehtaab Sawhney. On the real Davies’ conjecture. Annals of Probability, 49(6):3011– 3031, November 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-6/
On-the-real-Davies-conjecture/10.1214/21-AOP1522.full.

Kajino:2020:SEM

[KM20] Naotaka Kajino and Mathav Murugan. On singularity of energy measures for symmetric diffusions with full off-diagonal heat kernel estimates. Annals of Probability, 48(6):2920– 2951, November 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-6/
On-singularity-of-energy-measures-for-symmetric-diffusions-
with-full/10.1214/20-AOP1440.full.

Kosygina:2023:CNS

[KMP23] Elena Kosygina, Thomas Mountford, and Jonathon Peterson. Convergence and nonconvergence of scaled self-interacting random walks to Brownian motion perturbed at extrema. Annals of Probability, 51(5):1684–1728, September 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-5/Convergence-and-nonconvergence-
of-scaled-self-interacting-random-walks-to/10.1214/23-
AOP1629.full.

Komorowski:2020:FDL

[KOR20] Tomasz Komorowski, Stefano Olla, and Lenya Ryzhik. Fractional diffusion limit for a kinetic equation with an interface. Annals of Probability, 48(5):2290–2322, September 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-5/Fractional-diffusion-
limit-for-a-kinetic-equation-with-an-interface/10.1214/
20-AOP1423.full.

Kopel:2020:RMP

[KOV20] Phil Kopel, Sean O’Rourke, and Van Vu. Random matrix products: Universality and least singular values. Annals of Probability, 48(3):1372–1410, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1592359232.

Kenyon:2022:MTM

[KP22] Richard Kenyon and Cosmin Pohoata. The multinomial tiling model. Annals of Probability, 50(5):1986–2012, September 2022. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-50/issue-5/The-
multinomial-tiling-model/10.1214/22-AOP1575.full.

Krylov:2021:SED

[Kry21a] N. V. Krylov. On stochastic equations with drift in L_{d}. Annals of Probability, 49(5):2371–2398, September 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-5/On-stochastic-equations-
with-drift-in-Ld/10.1214/21-AOP1510.full.

Krylov:2021:SSI

[Kry21b] N. V. Krylov. On strong solutions of Itô’s equations with σ ∈ W_{d}^{1} and b ∈ L_{d}. Annals of Probability, 49(6):3142– 3167, November 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-6/
On-strong-solutions-of-It%c3%b4s-equations-with-%cf%83Wd1-
and-bLd/10.1214/21-AOP1525.full.

Krylov:2023:SSI

[Kry23] N. V. Krylov. On strong solutions of Itô’s equations with Dσ and b in Morrey classes containing L_{d}. Annals of Probability, 51(5): 1729–1751, September 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-5/
On-strong-solutions-of-It%c3%b4s-equations-with-D%cf%
83-and-b/10.1214/23-AOP1630.full.

Kosloff:2020:FIB

[KS20a] Zemer Kosloff and Terry Soo. Finitary isomorphisms of Brownian motions. Annals of Probability, 48(4):1966–1979, July 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-4/Finitary-isomorphisms-
of-Brownian-motions/10.1214/19-AOP1412.full.

Kulik:2020:WPS

[KS20b] Alexei Kulik and Michael Scheutzow. Well-posedness, stability and sensitivities for stochastic delay equations: a generalized coupling approach. Annals of Probability, 48(6):3041–3076, November 2020. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-48/issue-6/Well-
posedness-stability-and-sensitivities-for-stochastic-
delay-equations/10.1214/20-AOP1449.full.

Kozma:2023:LTR

[KS23] Gady Kozma and Wojciech Samotij. Lower tails via relative entropy. Annals of Probability, 51(2):665–698, March 2023.

CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-2/Lower-tails-via-relative-
entropy/10.1214/22-AOP1610.full.

Lacoin:2022:CLC

[Lac22] Hubert Lacoin. Convergence in law for complex Gaussian multiplicative chaos in phase III. Annals of Probability, 50(3): 950–983, May 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-3/
Convergence-in-law-for-complex-Gaussian-multiplicative-
chaos-in-phase/10.1214/21-AOP1551.full.

Lammers:2021:GHD

[Lam21] Piet Lammers. A generalisation of the honeycomb dimer model to higher dimensions. Annals of Probability, 49(2): 1033–1066, March 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-2/
A-generalisation-of-the-honeycomb-dimer-model-to-higher-
dimensions/10.1214/20-AOP1469.full.

LeGall:2022:GSR

[Le 22] Jean-François Le Gall. Geodesic stars in random geometry. Annals of Probability, 50(3):1013–1058, May 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-3/Geodesic-stars-in-random-
geometry/10.1214/21-AOP1553.full.

Lee:2021:CGR

[Lee21] James R. Lee. Conformal growth rates and spectral geometry on distributional limits of graphs. Annals of Probability, 49(6): 2671–2731, November 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-6/
Conformal-growth-rates-and-spectral-geometry-on-distributional-
limits-of/10.1214/20-AOP1480.full.

Liu:2022:MDT

[Liu22] Zhipeng Liu. Multipoint distribution of TASEP. Annals of Probability, 50(4):1255–1321, July 2022. CODEN APBYAE.

ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/journals/annals-of-probability/volume-
50/issue-4/Multipoint-distribution-of-TASEP/10.1214/21-
AOP1557.full.

Landon:2020:CTS

[LLM20] Benjamin Landon, Patrick Lopatto, and Jake Marcinek. Comparison theorem for some extremal eigenvalue statistics. Annals of Probability, 48(6):2894–2919, November 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-6/Comparison-theorem-for-
some-extremal-eigenvalue-statistics/10.1214/20-AOP1439.
full.

Lee:2020:HPB

[LMN20] Jong Jun Lee, Carl Mueller, and Eyal Neuman. Hitting probabilities of a Brownian flow with radial drift. Annals of Probability, 48(2):646–671, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1587542675.

Landim:2023:MBW

[LMS23] C. Landim, D. Marcondes, and I. Seo. Metastable behavior of weakly mixing Markov chains: The case of reversible, critical zero-range processes. Annals of Probability, 51(1):157– 227, January 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-1/
Metastable-behavior-of-weakly-mixing-Markov-chains--The-
case/10.1214/22-AOP1593.full.

Linker:2021:CPR

[LMSV21] Amitai Linker, Dieter Mitsche, Bruno Schapira, and Daniel Valesin. The contact process on random hyperbolic graphs: Metastability and critical exponents. Annals of Probability, 49 (3):1480–1514, May 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-3/
The-contact-process-on-random-hyperbolic-graphs--Metastability-
and/10.1214/20-AOP1489.full.

Lohr:2020:ACC

[LMW20] Wolfgang Löhr, Leonid Mytnik, and Anita Winter. The Aldous chain on cladograms in the diffusion limit. Annals of Probability, 48(5):2565–2590, September 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-5/The-Aldous-chain-on-
cladograms-in-the-diffusion-limit/10.1214/20-AOP1431.
full.

Landon:2023:KTF

[LNS23] Benjamin Landon, Christian Noack, and Philippe Sosoe. KPZtype fluctuation exponents for interacting diffusions in equilibrium. Annals of Probability, 51(3):1139–1191, May 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-3/KPZ-type-fluctuation-
exponents-for-interacting-diffusions-in-equilibrium/10.
1214/22-AOP1617.full.

Legrand:2022:STC

[LP22] Alexandre Legrand and Nicolas Pétrélis. Surface transition in the collapsed phase of a self-interacting walk adsorbed along a hard wall. Annals of Probability, 50(4): 1538–1588, July 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-4/
Surface-transition-in-the-collapsed-phase-of-a-self-interacting/
10.1214/22-AOP1567.full.

LeJan:2004:FCN

[LR04] Yves Le Jan and Olivier Raimond. Flows, coalescence and noise. Annals of Probability, 32(2):1247–1315, April 2004. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1084884851. See cor-

rection [LR20b].

LeGall:2020:GFP

[LR20a] Jean-François Le Gall and Armand Riera. Growth-fragmentation processes in Brownian motion indexed by the Brownian tree. Annals of Probability, 48(4):1742–1784, July 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-probability/volume-48/issue-4/Growth-fragmentation-
processes-in-Brownian-motion-indexed-by-the-Brownian/
10.1214/19-AOP1406.full.

LeJan:2020:FCN

[LR20b] Yves Le Jan and Olivier Raimond. Flows, coalescence and noise. A correction. Annals of Probability, 48(3):1592–1595, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/euclid.aop/

1592359239. See [LR04].

Lacker:2020:IMP

[LSZ20] Daniel Lacker, Mykhaylo Shkolnikov, and Jiacheng Zhang. Inverting the Markovian projection, with an application to local stochastic volatility models. Annals of Probability, 48(5):2189– 2211, September 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-5/
Inverting-the-Markovian-projection-with-an-application-
to-local-stochastic/10.1214/19-AOP1420.full.

Livshyts:2021:SSV

[LTV21] Galyna V. Livshyts, Konstantin Tikhomirov, and Roman Vershynin. The smallest singular value of inhomogeneous square random matrices. Annals of Probability, 49(3):1286–1309, May 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-3/The-smallest-
singular-value-of-inhomogeneous-square-random-matrices/
10.1214/20-AOP1481.full.

Lyons:2023:MCT

[LW23] Russell Lyons and Graham White. Monotonicity for continuoustime random walks. Annals of Probability, 51(3):1112–1138, May 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-51/issue-3/Monotonicity-
for-continuous-time-random-walks/10.1214/22-AOP1615.full.

Lyons:2013:DCM

[Lyo13] Russell Lyons. Distance covariance in metric spaces. Annals of Probability, 41(5):3284–3305, September 2013. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

http://projecteuclid.org/euclid.aop/1378991840. See er-

rata [Lyo18].

Lyons:2018:EDC

[Lyo18] Russell Lyons. Errata to “Distance covariance in metric spaces”. Annals of Probability, 46(4):2400–2405, July 2018. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1528876831. See

[Lyo13, Lyo21].

Lyons:2021:SED

[Lyo21] Russell Lyons. Second errata to “Distance covariance in metric spaces”. Annals of Probability, 49(5):2668–2670, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-5/Second-errata-
to-Distance-covariance-in-metric-spaces/10.1214/20-AOP1504.
full.

Meszaros:2020:LED

[Més20] András Mészaros. Limiting entropy of determinantal processes. Annals of Probability, 48(5):2615–2643, September 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-5/Limiting-entropy-of-
determinantal-processes/10.1214/20-AOP1435.full.

Mareche:2020:EAD

[MMT20] Laure Marêché, Fabio Martinelli, and Cristina Toninelli. Exact asymptotics for Duarte and supercritical rooted kinetically constrained models. Annals of Probability, 48(1):317–342, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL https://projecteuclid.org/euclid.
aop/1585123330.

Mousset:2020:PNB

[MNPS20] Frank Mousset, Andreas Noever, Konstantinos Panagiotou, and Wojciech Samotij. On the probability of nonexistence in binomial subsets. Annals of Probability, 48(1):493–525, January 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL https://projecteuclid.org/euclid.aop/1585123336.

Magazinov:2022:CIL

[MP22] Alexander Magazinov and Ron Peled. Concentration inequalities for log-concave distributions with applications to random surface fluctuations. Annals of Probability, 50(2):735–770, March 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-50/issue-2/Concentration-
inequalities-for-log-concave-distributions-with-applications-
to-random/10.1214/21-AOP1545.full.

Miclo:2022:DSS

[MPS22] Laurent Miclo, Pierre Patie, and Rohan Sarkar. Discrete selfsimilar and ergodic Markov chains. Annals of Probability, 50(6): 2085–2132, November 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-6/
Discrete-self-similar-and-ergodic-Markov-chains/10.1214/
22-AOP1577.full.

Muirhead:2023:PTP

[MRVKS23] Stephen Muirhead, Alejandro Rivera, Hugo Vanneuville, and Laurin Köhler-Schindler. The phase transition for planar Gaussian percolation models without FKG. Annals of Probability, 51(5): 1785–1829, September 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-5/
The-phase-transition-for-planar-Gaussian-percolation-
models-without-FKG/10.1214/23-AOP1633.full.

Miller:2021:LQG

[MS21] Jason Miller and Scott Sheffield. Liouville quantum gravity and the Brownian map II: Geodesics and continuity of the embedding. Annals of Probability, 49(6):2732–2829, November 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-6/Liouville-quantum-gravity-
and-the-Brownian-map-II--Geodesics/10.1214/21-AOP1506.
full.

Miller:2022:SCL

[MSW22] Jason Miller, Scott Sheffield, and Wendelin Werner. Simple conformal loop ensembles on Liouville quantum gravity. Annals of Probability, 50(3):905–949, May 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-3/Simple-conformal-loop-
ensembles-on-Liouville-quantum-gravity/10.1214/21-AOP1550.
full.

Mukherjee:2020:IPM

[MV20] Chiranjib Mukherjee and S. R. S. Varadhan. Identification of the Polaron measure in strong coupling and the Pekar variational formula. Annals of Probability, 48(5):2119– 2144, September 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-5/
Identification-of-the-Polaron-measure-in-strong-coupling-
and-the/10.1214/19-AOP1392.full.

Nam:2020:LDL

[Nam20] Kyeongsik Nam. Large deviations and localization of the microcanonical ensembles given by multiple constraints. Annals of Probability, 48(5):2525–2564, September 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-5/Large-deviations-and-
localization-of-the-microcanonical-ensembles-given-by/
10.1214/20-AOP1430.full.

Najnudel:2023:SCH

[NPS23] Joseph Najnudel, Elliot Paquette, and Nick Simm. Secular coefficients and the holomorphic multiplicative chaos. Annals of Probability, 51(4):1193–1248, July 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-4/Secular-coefficients-
and-the-holomorphic-multiplicative-chaos/10.1214/22-AOP1616.
full.

Nica:2020:SKE

[NQR20] Mihai Nica, Jeremy Quastel, and Daniel Remenik. Solution of the Kolmogorov equation for TASEP. Annals of Probability, 48(5): 2344–2358, September 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-48/issue-5/
Solution-of-the-Kolmogorov-equation-for-TASEP/10.1214/
20-AOP1425.full.

Nadtochiy:2020:MFS

[NS20] Sergey Nadtochiy and Mykhaylo Shkolnikov. Mean field systems on networks, with singular interaction through hitting times. Annals of Probability, 48(3):1520–1556, May 2020. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

https://projecteuclid.org/euclid.aop/1592359237.

Nutz:2023:SSP

[NW23] Marcel Nutz and Johannes Wiesel. Stability of Schrödinger potentials and convergence of Sinkhorn’s algorithm. Annals of Probability, 51(2):699–722, March 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/journals/annals-of-probability/volume-
51/issue-2/Stability-of-Schr%c3%b6dinger-potentials-and-
convergence-of-Sinkhorns-algorithm/10.1214/22-AOP1611.
full.

Pimentel:2022:IPK

[Pim22] Leandro P. R. Pimentel. Integration by parts and the KPZ two-point function. Annals of Probability, 50(5):1755– 1780, September 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-5/
Integration-by-parts-and-the-KPZ-two-point-function/10.
1214/22-AOP1564.full.

Perkowski:2021:RSB

[PR21] Nicolas Perkowski and Tommaso Rosati. A rough super-Brownian motion. Annals of Probability, 49(2):908–943, March 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-2/A-rough-super-
Brownian-motion/10.1214/20-AOP1464.full.

Priola:2021:ORR

[Pri21] Enrico Priola. An optimal regularity result for Kolmogorov equations and weak uniqueness for some critical SPDEs. Annals of Probability, 49(3):1310–1346, May 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-3/An-optimal-regularity-
result-for-Kolmogorov-equations-and-weak-uniqueness/10.
1214/20-AOP1482.full.

Pete:2022:FUS

[PT22a] Gábor Pete and Ádám Timár. The free uniform spanning forest is disconnected in some virtually free groups, depending on the generator set. Annals of Probability, 50(6):2218–2243, November 2022. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-50/issue-6/The-
Free-Uniform-Spanning-Forest-is-disconnected-in-some-
virtually/10.1214/22-AOP1581.full.

Pitman:2022:HSL

[PT22b] Jim Pitman and Wenpin Tang. Hidden symmetries and limit laws in the extreme order statistics of the Laplace random walk. Annals of Probability, 50(4):1647–1673, July 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-4/Hidden-symmetries-and-
limit-laws-in-the-extreme-order-statistics/10.1214/22-
AOP1572.full.

Paouris:2022:HLD

[PTV22] Grigoris Paouris, Konstantin Tikhomirov, and Petros Valettas. Hypercontractivity and lower deviation estimates in normed spaces. Annals of Probability, 50(2):688–734, March 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-2/Hypercontractivity-and-
lower-deviation-estimates-in-normed-spaces/10.1214/21-
AOP1543.full.

Raoufi:2020:TIG

[Rao20] Aran Raoufi. Translation-invariant Gibbs states of the Ising model: General setting. Annals of Probability, 48(2):760–777, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-

894X (electronic). URL https://projecteuclid.org/euclid.
aop/1587542679.

Riera:2022:IIB

[Rie22] Armand Riera. Isoperimetric inequalities in the Brownian plane. Annals of Probability, 50(5):2013–2055, September 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-50/issue-5/Isoperimetric-

inequalities-in-the-Brownian-plane/10.1214/22-AOP1576.
full.

Rossignol:2021:SLD

[Ros21] Raphaël Rossignol. Scaling limit of dynamical percolation on critical Erdős–Rényi random graphs. Annals of Probability, 49 (1):322–399, January 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-1/
Scaling-limit-of-dynamical-percolation-on-critical-Erd%
c5%91sR%c3%a9nyi-random-graphs/10.1214/20-AOP1472.full.

Rath:2021:FPB

[RST21] Balázs Ráth, Jan M. Swart, and Tamás Terpai. Frozen percolation on the binary tree is nonendogenous. Annals of Probability, 49(5): 2272–2316, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-5/
Frozen-percolation-on-the-binary-tree-is-nonendogenous/
10.1214/21-AOP1507.full.

Rockner:2021:DAF

[RX21] Michael Röckner and Longjie Xie. Diffusion approximation for fully coupled stochastic differential equations. Annals of Probability, 49(3):1205–1236, May 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://

projecteuclid.org/journals/annals-of-probability/volume-
49/issue-3/Diffusion-approximation-for-fully-coupled-
stochastic-differential-equations/10.1214/20-AOP1475.
full.

Remy:2020:DGM

[RZ20] Guillaume Remy and Tunan Zhu. The distribution of Gaussian multiplicative chaos on the unit interval. Annals of Probability, 48(2):872–915, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/euclid.aop/1587542682.

Salez:2023:UCE

[Sal23] Justin Salez. Universality of cutoff for exclusion with reservoirs. Annals of Probability, 51(2):478–494, March 2023. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-probability/volume-51/issue-2/Universality-of-cutoff-
for-exclusion-with-reservoirs/10.1214/22-AOP1600.full.

Schapira:2020:CRD

[Sch20a] Bruno Schapira. Capacity of the range in dimension. Annals of Probability, 48(6):2988–3040, November 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-6/Capacity-of-the-range-
in-dimension-5/10.1214/20-AOP1442.full.

Schmidt:2020:SPD

[Sch20b] Marius A. Schmidt. A simple proof of the DPRZ theorem for 2d cover times. Annals of Probability, 48(1):445–457, January 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/euclid.aop/
1585123334.

Schweiger:2020:MFD

[Sch20c] Florian Schweiger. The maximum of the four-dimensional membrane model. Annals of Probability, 48(2):714–741, March 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/euclid.aop/
1587542677.

Schiavo:2022:DFD

[Sch22]  Lorenzo Dello Schiavo.  The Dirichlet–Ferguson diffusion

on the space of probability measures over a closed Riemannian manifold. Annals of Probability, 50(2):591–648, March 2022. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-50/issue-2/The-
DirichletFerguson-diffusion-on-the-space-of-probability-
measures-over/10.1214/21-AOP1541.full.

Schmid:2023:MTT

[Sch23] Dominik Schmid. Mixing times for the TASEP in the maximal current phase. Annals of Probability, 51(4):1342–1379, July 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-51/issue-4/Mixing-times-for-the-
TASEP-in-the-maximal-current-phase/10.1214/22-AOP1620.
full.

Sellke:2022:CAR

[Sel22] Mark Sellke. Cutoff for the asymmetric riffle shuffle. Annals of Probability, 50(6):2244–2287, November 2022. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-6/Cutoff-for-the-asymmetric-
riffle-shuffle/10.1214/22-AOP1582.full.

Sjostrand:2023:MSL

[Sjö23] Jonas Sjöstrand. Monotone subsequences in locally uniform random permutations. Annals of Probability, 51(4): 1502–1547, July 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-4/
Monotone-subsequences-in-locally-uniform-random-permutations/
10.1214/23-AOP1624.full.

Spinka:2020:FCS

[Spi20a] Yinon Spinka. Finitary codings for spatial mixing Markov random fields. Annals of Probability, 48(3):1557–1591, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

https://projecteuclid.org/euclid.aop/1592359238.

Spinka:2020:FDP

[Spi20b] Yinon Spinka. Finitely dependent processes are finitary. Annals of Probability, 48(4):2088–2117, July 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-4/Finitely-dependent-processes-
are-finitary/10.1214/19-AOP1417.full.

Salins:2021:MEP

[SS21] Michael Salins and Konstantinos Spiliopoulos. Metastability and exit problems for systems of stochastic reaction–diffusion equations. Annals of Probability, 49(5):2317–2370, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-5/Metastability-
and-exit-problems-for-systems-of-stochastic-reactiondiffusion-
equations/10.1214/21-AOP1509.full.

Shen:2022:LLL

[SSZZ22] Hao Shen, Scott A. Smith, Rongchan Zhu, and Xiangchan Zhu. Large N limit of the O(N) linear sigma model via stochastic quantization. Annals of Probability, 50(1):131–202, January 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-50/issue-1/Large-N-limit-of-the-
ON-linear-sigma-model-via/10.1214/21-AOP1531.full.

Subag:2023:TAM

[Sub23] Eliran Subag. TAP approach for multispecies spherical spin glasses II: The free energy of the pure models. Annals of Probability, 51 (3):1004–1024, May 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-3/
TAP-approach-for-multispecies-spherical-spin-glasses-
II--The/10.1214/22-AOP1605.full.

Sarkar:2021:BAC

[SV21] Sourav Sarkar and Bálint Virág. Brownian absolute continuity of the KPZ fixed point with arbitrary initial condition. Annals of Probability, 49(4):1718–1737, July 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-4/Brownian-absolute-continuity-
of-the-KPZ-fixed-point-with-arbitrary/10.1214/20-AOP1491.
full.

Saksman:2020:RZF

[SW20] Eero Saksman and Christian Webb. The Riemann zeta function and Gaussian multiplicative chaos: Statistics on the critical line. Annals of Probability, 48(6):2680–2754, November 2020. CODEN APBYAE. ISSN 0091-1798 (print),

2168-894X (electronic). URL https://projecteuclid.org/
journals/annals-of-probability/volume-48/issue-6/The-
Riemann-zeta-function-and-Gaussian-multiplicative-chaos-
-Statistics/10.1214/20-AOP1433.full.

Sabot:2020:HTI

[SZ20] Christophe Sabot and Xiaolin Zeng. Hitting times of interacting drifted Brownian motions and the vertex reinforced jump process. Annals of Probability, 48(3):1057–1085, May 2020. CODEN AP-BYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL

https://projecteuclid.org/euclid.aop/1592359222.

Sly:2022:SDV

[SZ22] Allan Sly and Lingfu Zhang. Stationary distributions for the voter model in d ≥ 3 are factors of IID. Annals of Probability, 50(4):1589–1609, July 2022. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-50/issue-4/
Stationary-distributions-for-the-voter-model-in-d3-are-
factors/10.1214/22-AOP1568.full.

Teyssier:2020:LPR

[Tey20] Lucas Teyssier. Limit profile for random transpositions. Annals of Probability, 48(5):2323–2343, September 2020. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-48/issue-5/Limit-profile-for-random-
transpositions/10.1214/20-AOP1424.full.

Thevenin:2021:GRF

[Thé21] Paul Thévenin. A geometric representation of fragmentation processes on stable trees. Annals of Probability, 49(5):2416– 2476, September 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-49/issue-5/
A-geometric-representation-of-fragmentation-processes-
on-stable-trees/10.1214/21-AOP1512.full.

Thoma:2023:TSL

[Tho23] Eric Thoma. Thermodynamic and scaling limits of the non-Gaussian membrane model. Annals of Probability, 51(2): 626–664, March 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-2/
Thermodynamic-and-scaling-limits-of-the-non-Gaussian-
membrane-model/10.1214/22-AOP1609.full.

Timar:2021:NFE

[Tim21] Ádám Timár. A nonamenable “factor” of a Euclidean space. Annals of Probability, 49(3):1427–1449, May 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-probability/volume-49/issue-3/A-nonamenable-factor-
of-a-Euclidean-space/10.1214/20-AOP1485.full.

Tolomeo:2021:GWP

[Tol21] Leonardo Tolomeo. Global well posedness of the two-dimensional stochastic nonlinear wave equation on an unbounded domain. Annals of Probability, 49(3):1402–1426, May 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-probability/volume-49/issue-3/Global-well-posedness-
of-the-two-dimensional-stochastic-nonlinear-wave/10.1214/
20-AOP1484.full.

Vanneuville:2021:ASS

[Van21] Hugo Vanneuville. The annealed spectral sample of Voronoi percolation. Annals of Probability, 49(3):1554–1606, May 2021. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/journals/
annals-of-probability/volume-49/issue-3/The-annealed-
spectral-sample-of-Voronoi-percolation/10.1214/20-AOP1494.
full.

Valko:2020:OLC

[VV20] Benedek Valkó and Bálint Virág. Operator limit of the circular beta ensemble. Annals of Probability, 48(3):1286–1316, May 2020. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X

(electronic). URL https://projecteuclid.org/euclid.aop/
1592359229.

Xiao:2023:LDE

[XGL23] Hui Xiao, Ion Grama, and Quansheng Liu. Large deviation expansions for the coefficients of random walks on the general linear group. Annals of Probability, 51(4):1380– 1420, July 2023. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic). URL https://projecteuclid.

org/journals/annals-of-probability/volume-51/issue-4/
Large-deviation-expansions-for-the-coefficients-of-random-
walks-on/10.1214/23-AOP1621.full.

Xu:2021:STI

[Xu21] Changji Xu. Sharp threshold for the Ising perceptron model. Annals of Probability, 49(5):2399–2415, September 2021. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (elec-

tronic). URL https://projecteuclid.org/journals/annals-

REFERENCES 72

of-probability/volume-49/issue-5/Sharp-threshold-for-
the-Ising-perceptron-model/10.1214/21-AOP1511.full.